### v1

# LASLiN: A Learning-Augmented Peer-to-Peer Network

Abstract
We introduce a learning-augmented peer-to-peer (P2P) network design that leverages the predictions of traffic patterns to optimize the network’s topology. While keeping formal guarantees on the standard P2P metrics (routing path length, maximum degree), we optimize the network in a demand-aware manner and minimize the path lengths weighted by the peer-to-peer communication demands. Our protocol is learning-augmented, meaning that each node receives an individual, possibly inaccurate prediction about the future traffic patterns, with the goal of improving the network’s performances. In this paper, we strike a trade-off between significantly improved performances when the predictions are correct (consistency) and polylogarithmic performances when the predictions are arbitrary (robustness). Our protocol builds upon a generalization of skip lists to networks.
We have two main contributions. First, we consider the centralized setting and show that the problem of constructing an optimum static skip list network (SLN) is solvable in polynomial time and can be computed via dynamic programming. This problem is the natural demand-aware extension of the optimal skip list problem; in particular, we use it to generate the predictions of our learning-augmented protocol.
Second, we introduce the Uniform P2P protocol which generalizes skip list networks (SLN) by allowing the nodes to have any real-valued height instead of discrete ones. We show that Uniform achieves state-of-the-art performances: logarithmic routing and maximum degree, both with high probability. We then use Uniform to build a learning-augmented P2P protocol in order to incorporate demand-awareness, leading to our main contribution, LASLiN. We prove that the performances of LASLiN are consistent with those of an optimum static SLN with correct predictions, and are at most a logarithmic factor off the state-of-the-art P2P protocols if the predictions are arbitrary wrong. For the special case of highly sparse demands, we show that LASLiN achieves an exponential improvement in case of correct predictions, and asymptotically matches the performances of state-of-theart, demand-oblivious P2P protocols in case of adversarial predictions.
keywords:
learning-augmented algorithms, peer-to-peer networks, skip list graphs
1 Introduction
Peer-to-peer (P2P) protocols are one of the key technological advancements that allowed decentralized and distributed systems to scale. A vast body of work has been devoted to optimize the efficiency of these protocols [19, 27, 3, 16, 17]. Traditionally, the efficiency of a P2P protocol is estimated with two key metrics which are the routing path length between a pair of nodes and the maximum degree of each node in the network. State-of-the-art P2P protocols achieve both a logarithmic routing path length and a logarithmic maximum degree [29, 13].
Up until recently, most work on P2P protocols focused on demand-oblivious worst-case scenarios, where no prior information about the possible future demand is known. However, a new line of work studies demand-aware networks [28, 5, 12] which can adapt their connectivity properties to better fit the current traffic. Noting that most networking traffics are sparse and partly predictable [4], the idea is to detect the few highly active nodes and move them topologically closer to reduce the overall delays. The model includes a demand matrix which indicates the amount of traffic between any two nodes in the network; A goal of a demand-aware protocol is then to minimize the total weighted path length [28].
This paper presents a P2P network design where each node can adapt its set of neighbours to better fit the future traffic demands. We do not assume that the demand matrix is known by the nodes; Rather, we stick with the distributed nature of P2P protocols and assume that each node has access to a local prediction about the future demands — the prediction may only consist of partial information about the future traffic, an integer number in the case of this work. In our model, a node may obtain a new prediction at any time to use at its own discretion, independently from the other nodes. Crucially, the predictions are untrusted and we evaluate our P2P protocol in the learning-augmented setting with two metrics: consistency and robustness [20]. The consistency of the algorithm is its performance when the predictions are always perfect, its robustness is its performance when the predictions are worst possible.
1.1 Our Setup
P2P Networks
A P2P protocol is an algorithm distributed over multiple nodes that provides routing facilities between its member nodes. On top of a physical, static network, the P2P protocol keeps track of local routing information stored inside each node.
This paper studies P2P network protocols in which each node $u$ holds a height variable $h_{u}$. In addition, we assume that each node holds a unique identifier (id) which defines a total order at any time: for any two different nodes $u$ and $v$, either $u<v$ or $v<u$. A Skip List Network (SLN[21] is a P2P design based on the well-known skip list data structure [25] which establishes the routing informations solely based on the heights and ids of the nodes. In a SLN, two nodes $u$ and $v$ form an edge, i.e., store each other’s id, if and only if there is no node $x$ such that $u<x<v$ and $h_{x}\geq\min(h_{u},h_{v})$ — see Figure 1.
P2P Networks Constraints
Following the overall consensus [13, 23] on P2P networks, we restrict our focus to P2P protocols with the following properties: (1) any node should be able to join or leave the network at any time, (2) the routing should be local, i.e., each new hop through the network is decided solely by one node, and (3) the routing should be oblivious, i.e., the id of the destination node is sufficient to route the message.
Stated the above requirements, we want the network to still work at scale when the number of nodes in the network becomes large. Noting $n$ the number of nodes at the considered time, we aim at networks with:
  • small routing path lengths. Given any two nodes $u$ and $v$, at all time, the routing path length between $u$ and $v$ is ideally logarithmic in $n$ with high probability.
  • small degrees. At all time, the number of known nodes for any node, i.e., the maximum degree, is ideally logarithmic in $n$ with high probability.
All the P2P protocols described of this paper respect those two criteria. To evaluate the performances of our P2P protocols, we compare them to idealized P2P protocols with the restriction that they need to have polylogarithmic path length and maximum degree (see Definition 2.4).
Traffic Demand
At a given time, we assume that the traffic demand between the nodes is modelled with a demand matrix $W\in M_{n}(\mathbb{R})$, where $W(u,v)$ is the frequency of communications from $u$ to $v$ [28].
Learning-Augmented Protocols
This paper introduces and studies learning-augmented P2P protocols. In the general case, a learning-augmented algorithm is an algorithm that takes an additional input, the prediction, to improve its performances. Crucially, the prediction is untrusted. This means that, with regard to the other inputs of the algorithm, the prediction may be perfect (all things being equal, no other prediction results in a better performance), adversarial, or anything in between.
In our design of learning-augmented P2P protocols, we assume that each node can obtain a prediction at any time, independently from one another. The origin of the prediction is not the focus of this paper; it may come from a centralized server handing out traffic predictions to the nodes, or it may be locally computed by each node based on its own traffic history. In this work, the input prediction is an integer interpreted as the height of the node in the SLN.
1.2 Contributions and Technical Novelty
The contributions of this paper are threefold. First, we formulate the Optimum Static Skip List Network (OSSLN) problem and prove it can be solved in polynomial time, with an algorithm based on dynamic programming (Theorem 3.3). This result extends the previous work of [22] to the case with pairwise demands in line with recent demand-aware results [28]. Later on, this algorithm demonstrates that the task of finding optimal predictions for our learning-augmented P2P protocol is also polynomial.
012312345678
Figure 1: A representation of a Skip List Network (SLN) with eight nodes. The nodes have ids ranging from $1$ to $8$. The height of a rectangle equals the height of its corresponding node; Notice that all heights have integer values as this is an SLN. The double-arrowed, horizontal lines are the edges of the SLN.
0112345678
Figure 2: A continuous SLN with real-valued heights in $(0,1)$.
0123412345678
Figure 3: A Continuous SLN obtained by adding node-by-node the heights of the SLN in Figure 1 and the heights of the Continuous SLN in Figure 2. This is a possible outcome of LASLiN where the predicted heights are those of Figure 1 and the uniform, random heights are those of Figure 2.
01212345678
Figure 4: A SLN with all equal heights, effectively turning the SLN into a simple chained list. This SLN has a linear routing path length. This worst-case performance happens when all the nodes get the same predicted height.
01212345678
Figure 5: The addition of continuous, random heights restores the good performances.
As our second contribution, we present a new P2P design called Continuous Skip List Network (Continuous SLN, Definition 4.1), which generalizes Skip List Networks [6] (SLN, Definition 2.1) and skip lists [25]. We show that a SLN with real-valued heights enables a functioning P2P protocol in that we implement the three basic primitives route, join and leave with state-of-the-art, logarithmic message complexity. We note that all known P2P designs that are based on skip lists [16, 14, 3] historically perceive it as chained lists stacked on each other and explicitly use that discrete nature in the protocol implementation, we show that this is not necessary. We then present a Continuous SLN protocol called Uniform that draws its heights uniformly at random in the interval $(0,1)$ and show that Uniform matches the performances of state-of-the-art P2P protocols with both routing path lengths and maximum degrees lower than $O(\log n)$ with high probability.
As our third contribution, we present a learning-augmented P2P protocol, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform protocol with a predicted SLN to strike consistency and robustness guarantees (Definition 2.8). For each node, LASLiN takes as prediction an integer between $1$ and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $(0,1)$ uniformly at random and adds it to the integer heights to obtain its final heights as a Continuous SLN. We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $\ln n$ factor off a demand-oblivious P2P protocol with arbitrary predictions ($O(\log^{2}n)$-robust). Finally we show that, in the special case of a very sparse demand matrix (see Definition 6.5), the version of LASLiN with binary predictions always outperforms any demand-oblivious P2P protocol with high probability: LASLiN obtains an exponentially better cost with correct predictions while it asymptotically matches the cost of the state-of-the-art P2P protocols for arbitrary predictions (see Theorem 6.6). We provide simulations to assess the performance of LASLiN on Zipfian demand matrices.
2 Model and Metrics
In this section, we formally define our model and the metrics we use to evaluate our P2P protocols.
Definition 2.1 (SLN).
A Skip List Network (SLN) $\mathcal{N}$ is a triple $(V,h,E)$ where:
  • $V$ is a totally ordered set of node ids;
  • $h$ is a height function $h:V\longrightarrow\mathbb{N}^{+}$;
  • $E=\{\{u,v\}\;|\;\forall x\in V\;s.t.\;u<x<v,\;h_{x}<\min(h_{u},h_{v})\}$.
Definition 2.2 (Neighbor and Degree).
We say that two nodes $u$ and $v$ are neighbors in a SLN $\mathcal{N}=(V,h,E)$ if $\{u,v\}\in E$. The degree of a node $u$ is its number of neighbors.
We redefine the greedy, local routing path proper to SLN.
Definition 2.3 (Routing Path).
Let $\mathcal{N}=(V,h,E)$ be a SLN and $u,v\in V$ be two different nodes. The routing path from $u$ to $v$ in $\mathcal{N}$ is the sequence of edges of $E$ $s=e_{1},e_{2}\dots e_{k}$ such that: (1) $u\in e_{1}$, (2) $v\in e_{k}$, and (3) $\forall i\in[k]$, if $u<v$ then $e_{i}=\{x,y\}$ where $y$ is the largest node neighbor of $x$ such that $x<y\leq v$, otherwise if $v<u$ then $e_{i}=\{x,y\}$ where $y$ is the smallest node neighbor of $x$ such that $v\leq y<x$.
More intuitively, the routing path goes to the next neighbor that is the closest to the destination (in terms of ids) without “flying over” the destination.
Definition 2.4 (Ideal P2P Protocol).
We say that a (potentially randomized) P2P protocol is Ideal if, at all times, the routing path length between two nodes and the maximum degree are polylogarithmic on $n$ with high probability, where $n$ is the number of nodes at the considered time. A quantity is polylogarithmic when there exists positive constants $c$ and $d$ such that the quantity is smaller than $d\cdot\log^{c}n$.
Definition 2.5 (Cost).
Let $\mathcal{N}=(V,h,E)$ be an SLN and $W$ be a demand matrix. The cost function is:
$\text{Cost}(\mathcal{N},W)=\sum_{u,v\in V}W(u,v)\cdot d_{\mathcal{N}}(u,v),$
where $d_{\mathcal{N}}(u,v)$ is the routing path length between $u$ and $v$ in $\mathcal{N}$.
Definition 2.6 (Optimum Static Skip List Network (OSSLN) problem).
Given a set of node ids $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\longrightarrow\mathbb{N}^{+}$ that minimizes the cost of the SLN $\mathcal{N}=(V,h,E)$.
Definition 2.7 (Learning-augmented P2P protocol).
A learning-augmented P2P protocol ALG is a P2P protocol whose initialization routine takes an additional parameter $p$, the prediction.
In our model, a node that wishes to update its prediction as it already joined the network simply leaves, re-intializes and joins again.
Definition 2.8 (Consistency and Robustness).
We say that a learning-augmented SLN protocol is
  • $\alpha$-consistent if, for all demand matrices, there exists an assignment of predictions to the nodes such that the expected cost of the protocol is at most $\alpha$ times the cost of an ideal optimum static SLN.
  • $\beta$-robust if, for all demand matrices and for all assignment of predictions, the expected cost of the protocol is at most $\beta$ times the cost of an ideal optimum static SLN.
3 Optimum Static Skip List Network (OSSLN)
Given a set of node ids $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\longrightarrow\mathbb{N}^{+}$ that minimizes the total communication cost (as defined in (2.5)) of the SLN $\mathcal{N}=(V,h,E)$ (as defined in (2.1)):
$OSSLN(V,W)=\min_{h}\;Cost(\mathcal{N},W).$ (1)
In [28], the authors presented an algorithm based on dynamic programming to find the Optimum static distributed Binary Search Tree (OBST) in polynomial time ($O(n^{3})$). This algorithm leverages a crucial property: once a vertex $v$ is chosen as the root of a tree induced by the vertices in a given interval, the subproblems induced by the vertex id intervals to the left and to the right of $v$ can be solved independently. The OSSLN problem exhibits a similar structure, but with crucial difference.
If we define a pivot vertex $x\in[l,r]$, $1\leq l\leq r\leq n,$ such that $h_{x}=\max_{w\in[l,r]}h_{w}$, $x$ also partitions the OSSLN problem into two subproblems (on the intervals $[l,v)$ and $(v,r]$). However, a critical distinction arises when $h_{x}<\max_{w\in[l,r]}h_{w}$, in which case $x$ will not be a part of all the paths between vertex pairs $(u,v)$ $u\in[l,x),v\in(x,r]$. Specifically, for a vertex $x\in[l+1,r-1]$ such that $h_{l},h_{r}>h_{x}$, any path between vertices $(u,v)$ where $u\leq l$ and $v\geq r$ does not include $x$. This changes the optimal substructure of the problem, requiring a more sophisticated dynamic programming approach.
Given a SLN $\mathcal{N}$, we define a path $P_{h}(l,r)$ to be the set of vertices in the routing path between the vertices $l$ and $r$ (as defined in (2.3)), excluding $l$, i.e., $l\notin P_{h}(l,r)$. This is a crucial simplification that allows the problem to be decomposed. Under this assumption, we have $d(l,r)=|P_{h}(l,r)|$. W.l.o.g., we assume that $P_{h}(l,r)=P_{h}(r,l)$.
Next, we define $L(x,h)=\min\{i\leq x\;|\;h_{j}\leq h_{x},\;\forall j\in[i,x]\}$, i.e., the vertex of minimum id between the set of all (consecutive) vertices of height lower or equal to $h_{x}$, to the left of $x$, and, analogously, $R(x,h)=\max\{i\geq x\;|\;h_{j}\leq h_{x},\;\forall j\in[x,i]\}$, i.e., the vertex with maximum id between the set of all vertices of height lower or equal to $h_{x}$, to the right of $x$. A vertex $x\in V$ belongs to a path $P_{h}(u,v)$ if and only if there does not exist any pair of vertices $(u^{\prime},v^{\prime})$ such that $u\leq u^{\prime}<x$, $x<v^{\prime}\leq v$, and (both) $h_{u^{\prime}},h_{v^{\prime}}>h_{x}$.
Finally, given an arbitrary $x\in V$, we define $P_{x}(h)$ as the set of all pairs $(u,v)\in V\times V$, such that $x\in P_{h}(u,v)$.
Lemma 3.1.
If $h$ is a solution for the $OSSLN(V,W)$, then $\forall x\in V$ we have that:
$P_{x}(h)$ $=$ $\{(u,v)\mid u<x\;\wedge\;v\in[x,R(x,h)]\}\;\cup$
$\{(u,v)\mid u\in[L(x,h),x)]\;\wedge\;v>R(x,h)\}$
Proof 3.2.
If $u<x$ and $x\leq v\leq R(x,h)$, then, by definition of $R(x,h)$, $h_{x}\geq h_{v}\geq\min(h_{u},h_{v})$. Since $x$ lies between $u$ and $v$ and there are no edges “above" $h_{x}$ in the interval $[u,v]$ (i.e., $\nexists(w,z)\in E$, s.t. $h_{w}>h_{x}$ and $h_{z}>h_{x}$, $w,z]\subseteq[u,v]$), $x$ must be a part of $P_{h}(u,v)$.
Similarly, if $L(x,h)\leq u<x$ and $v>R(x,h)$, then, again by definition of $L(x,h)$, $h_{x}\geq h_{u}\geq\min(h_{u},h_{v})$. Since $x$ lies between $u$ and $v$ and there are no edges "above" $h_{x}$ in the interval $[u,v]$, $x$ must be a part of $P_{h}(u,v)$
It remains to show that no other pair $(u,v)$ belongs to $P_{x}(h)$. If $u\geq x$, then by definition of $P_{h}(u,v)$, $x\notin P(u,v)$. Otherwise, suppose $u<L(x,h)$ and $v>R(x,h)$. Let $u^{\prime}=L(x,h)-1$ and $v^{\prime}=R(x,h)-1$, by definitions of $L(x,h)$ and $R(x,h)$, we have that $h_{x}<h_{u^{\prime}}$ and $h_{x}<h_{v^{\prime}}$. Since $u\leq u^{\prime}<x<v^{\prime}\leq v$, we have that:
$|P_{h}(u,u^{\prime})\cup\{v^{\prime}\}\cup P_{h}(v^{\prime},v)|<|P_{h}(u,x)\cup\{x\}\cup P_{h}(x,v)|.$
Because $P_{h}(u,v)$ is the shortest path from $u$ to $v$, it follows that $x\notin P_{h}(u,v)$.
Since $d(u,v)=|P_{h}(u,v)|$, we can now rewrite the total cost of an SLN $\mathcal{N}=(V,h,E)$ as:
$\text{Cost}(\mathcal{N},W)$ $=$ $\sum_{u,v\in V}W(u,v)\cdot d(u,v)$
$=\sum_{u,v\in V}\sum_{x\in P_{h}(u,v)}W(u,v)$ $=$ $\sum_{x\in V}\sum_{(u,v)\in P_{x}(h)}W(u,v).$
This means that the total cost is equal to the sum of contributions of all vertices $x\in V$, where the contribution of a vertex $x$ to the total cost is defined as:
$\sum_{(u,v)\in P_{x}(h)}W(u,v)$ (3)
OSSLN with Bounded Height: In the OSSLN with bounded height, the height assignment is of the form $h:V\longrightarrow[1,h_{\max}]$, where $h_{\max}$ is a parameter, typically bounded by $O(\log{n})$ or $O(1)$.
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$. Let $nl$, $1\leq nl\leq l$, denote the leftmost vertex id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let $nr$, $r\leq nr\leq n$, denote the rightmost vertex id, such that $\forall v\in[r,nr],\,h_{v}\leq h$. Finally, let $\text{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all vertices $v\in[l,r]$.
Theorem 3.3.
If $x\in[l,r]$ is a candidate pivot vertex with height $h\in[1,h_{\max}]$, then the following recurrence relation holds $\forall(nl,l,x,r,nr)$, $1\leq nl\leq l\leq x\leq r\leq nr\leq n$:
$\text{OSSLN}(l,r,nl,nr,h,W)=\min\Bigg{\{}\text{OSSLN}(l,r,l,r,h-1,W),$
$\min_{x\in[l,r]}\Bigg{[}\text{OSSLN}(l,x-1,nl,nr,h,W)+$
$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\Bigg{]}\Bigg{\}}.$
Proof 3.4.
The proof is by strong induction on the size of the interval $l,r]\subseteq[1,n]$:
Note that, by definition, $nl=L(x,h)$ and $nr=R(x,h)$ for any possible $h$. Furthermore, by construction, for an interval induced by $l$, $r$, nlnl, nrnr with bounded height $h$, assigning a pivot vertex with height equal to h1h-1, we have $nl^{\prime}=l$ and $nr^{\prime}=r$.
Base case $(l=r)$: In this case, we have that $(x=l)$ and the set of paths $x$ belongs to is given by Lemma 3.1. By setting $nl=L(x,h)$ and $nr=R(x,h)$ in the proof, the vertex contribution to the total cost is given by:
$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)$ (4)
Induction step: For $(h=1)$, we can disregard the first half of the recursion, where we assign the pivot vertex a height lower than $h$. The total contribution of all vertices to the cost becomes:
$\sum_{x\in[l,r]}\left(\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\right),$
which, from the base case, is precisely what the second half of the recurrence equation will return for any choice of the pivot vertex.
Let $1\leq l<r\leq n$ and $h\in[1,h_{\max}]$. Assume that the recurrence equation is true for all $h^{\prime},l^{\prime},r^{\prime}$ such that $(h^{\prime}<h)\leavevmode\nobreak\ \wedge\leavevmode\nobreak\ (r^{\prime}-l^{\prime})<(r-l)$. From previous step also assume that the equation is true for $(r^{\prime}-l^{\prime})=(r-l)$ with height $(h^{\prime}<h)$.
Since, when assigning the pivot vertex with height lower than $h$, by construction, we have that $(nl=l)$ and $(nr=r$). The minimum cost for the solution for any $nl$ and $nr$ is the minimum between keeping the height of the pivot, $(h_{x}=h)$, or decreasing it, $h_{x}\leq(h-1)$, which is given by the minimum between:
$\min_{x\in[l,r]}\Bigg{[}\text{OSSLN}(l,x-1,nl,nr,h,W)+$
$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\Bigg{]}$
and OSSLN(l,r,l,r,h1,W)OSSLN(l,r,l,r,h-1,W), which is precisely what we want to prove.
If we pre-compute an auxiliary matrix of aggregate weights ($W^{\prime}(x,y)=\sum_{z\leq y}W(x,z),\forall(x,y),1\leq x\leq y\leq n$) in time $O(n^{3})$, then the above dynamic programming algorithm has complexity $\mathcal{O}(n^{5}\cdot h_{\max})$.
OSSLN with Unbounded height (OSSLNU): We now present a dynamic programming algorithm for the OSSLN problem with unbounded heights, i.e., when the height assignment is of the form $h:V\longrightarrow[1,n]$. We show that, in this particular case, the algorithm’s complexity can be reduced from $\mathcal{O}(n^{5}\cdot h_{\max})$ to $\mathcal{O}(n^{3})$.
A crucial observation is that, for any solution to the OSSLNU problem, there exists a solution of cost less than or equal to that of the former, in which, $\forall[l,r],1\leq l\leq r\leq n$, there is a unique pivot vertex, i.e., a unique vertex whose height is the maximum height in that interval.
Lemma 3.5.
Let $h$ be an optimum solution to $OSSLNU(V,W)$. Then, there must exist a solution $h^{\prime}$ such that $Cost((V,h^{\prime},E^{\prime}),W)\leq Cost((V,h,E),W)$, where, for any interval $l,r],1\leq l\leq r\leq n,$ there is exactly one vertex $x\in[l,r]$ with height $h^{\prime}_{x}=\max_{w\in[l,r]}h^{\prime}_{w}$.
Proof 3.6.
Suppose, for contradiction, that any possible optimal solution contains at least one interval $[l,r]$ with at least two vertices $x,y\in[l,r]$ with $h_{x}=h_{y}=\max(h_{l,r})$ and $x\neq y$.
Define $h^{\prime}$ as follows for any vertex $v\in V$:
$h^{\prime}_{v}=\begin{cases}h_{v}+1&\text{if }v=x\text{ or }h_{v}>h_{x}\\ h_{v}&\text{otherwise}\end{cases}$
Note that, while $h^{\prime}_{x}>h^{\prime}_{y}$, this height assignment is order preserving. If $h_{a}>h_{b}$ then $h^{\prime}_{a}>h^{\prime}_{b}$.
If $x\notin P_{h}(u,v)$, since we preserved the order between $x$ and all vertices taller than $x$, $x$ cannot appear as an obstacle between any such vertices. Analogously, since we preserved the order between all vertices with at most the same height as $x$ there can’t be any new edges between vertices with at most the same height as $x$, so the path is unmodified.
If $x\in P_{h}(u,v)$, we know that:
$P_{h}(u,v)=P_{h}(u,x)\;\cup\{x\}\;\cup P_{h}(x,v).$
From our definition of path, we know that $x\notin P_{h}(x,v)$, and therefore $P_{h}(x,v)=P_{h^{\prime}}(x,v)$. Now, let’s consider a vertex $z$, such that $u\leq z<x$ and $h_{z}>h_{x}$. If such a vertex $z$ does not exist, then we can infer that $P_{h}=P_{h^{\prime}}$. Furthermore, if there does exist such a vertex $z$, then:
$|P_{h^{\prime}}|=|P_{h}(u,z)\;\cup\{x\}\;\cup P_{h}(x,v)|\leq|P_{h}|.$
Since $\text{Cost}((V,h,E),W)=\sum_{(u,v)\in V}W(u,v)\cdot d(u,v)$ and $d(u,v)=|P_{h}(u,v)|$, and any path in $h^{\prime}$ is of at most the same size as in $h$, then, $\text{Cost}((V,h^{\prime},E^{\prime}),W)\leq\text{Cost}((V,h,E),W)$. Hence, the lemma holds.
In OSSLNU, because we are only interested in the relative order of vertex heights, rather than in their absolute values, we have that $(nl=l)$ and $(nr=r)$, $\forall[l,r],1\leq l\leq r\leq n$, and the height parameter $h$ becomes unnecessary. This greatly simplifies the recurrence relation, reducing the complexity to $\mathcal{O}(n^{3})$.
Theorem 3.7.
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$. Let $\text{OSSLNU}(l,r,W)$ denote the minimum total cost contributed by all vertices $v\in[l,r]$. Let $x\in[l,r]$ be a candidate pivot vertex. Define $l$ as the leftmost index such that $\forall x,\in[l,r],\forall v\in[l,x],\,h_{v}\leq h_{x}$, and similarly define $r$ as the rightmost index such that $\forall v\in[x,r],\,h_{v}\leq h_{x}$. Then:
$\text{OSSNU}(l,r,W)=\min_{x\in[l,r]}\Big{(}\text{OSSNU}(l,x-1,W)+$
$\text{OSSNU}(x+1,r,W)+$
$\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)\Big{)}.$ (5)
Proof 3.8.
The proof is by induction on the size of the interval $[l,r]$.
Note that, by Lemma 3.5, once a pivot vertex $x\in[l,r]$ is chosen as the highest vertex within the interval $[l,r]$, we can assume, without loss of generality, that all vertices in the optimal solutions to the left subinterval [l,x1][l,x-1] and to the right subinterval $[x+1,r]$ have heights strictly lower than $h_{x}$.
Therefore, for a solution $h$ to $\text{OSSLNU}(l,r,W)$, it follows that $l=L(x,h)$ and $r=R(x,h)$.
Base Case $(l=r)$: In this case, the interval contains a single vertex $x=l=r$. In this case, the cost reduces to the contribution of $x$ to the solution cost, which, from Lemma 3.1, and $L(x,h)=l$ and $R(x,h)=r$ as we demonstrated above, it follows that this expression corresponds exactly to:
$\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)$ (6)
Inductive Step: Suppose the recurrence holds for all subintervals of $[l,r]$ of size less than (rl+1)(r-l+1). For an interval $[l,r]$ of size greater than one, we consider a candidate pivot vertex $x\in[l,r]$.
The total cost of placing $x$ as pivot of the interval $[l,r]$ equals to the sum of:
  • The optimal cost for the left subinterval [l,x1][l,x-1],
  • The optimal cost for the right subinterval $[x+1,r]$,
  • The cost of all paths that include $x$, which by Lemma 3.1 is given by the Eq.(6).
Since we minimize this total over all pivot choices $x\in[l,r]$, the recurrence holds.
4 Continuous SLN
This section introduces Continuous SLN, a generalization of SLN.
Definition 4.1 (Continuous SLN).
A Continuous Skip List Network (Continuous SLN) $\mathcal{N}=(V,h,E)$ is an SLN (Definition 2.1) except the height function $h$ is defined over the real numbers, i.e., $h:V\longrightarrow\mathbb{R}$.
We prove that continuous skip graphs support a functioning P2P network by implementing the three primitives: route, join and leave.
Internal Variables.
A node $u$ holds the variables:
  • $u$: the unique id of the node
  • $h_{u}$: the height in $\mathbb{R}$ of node $u$
  • $N_{u}$: the set of all neighbors of $u$
  • $D_{u}$: dictionary mapping a neighbor $w\in N_{u}$ to its height
  • $\text{anchor}_{u}$: the node with an id closest to $u$, either the successor or predecessor of $u$.
Route.
The route algorithm (Algorithm 1) takes the destination node id $v$ and hops from node to node until it finds the destination node, returning $(v,True)$ or reaches a dead end in some node $w\neq v$, return $(w,False)$.
1 // code running at node $u$
2 if $u$ = $v$ then
3   return ($u$, True)
4   else if $u<v$ then
5     $S\leftarrow\{x\in N_{u}|u<x\leq v\}$
6     if $S=\emptyset$ then
7       return ($u$, False)
8       else
9         $w\leftarrow\max S$
10         return $w.\text{route}(v)$
11
12         else
13           $S\leftarrow\{x\in N_{u}|v\leq x<u\}$
14           if $S=\emptyset$ then
15             return ($u$, False)
16             else
17               $w\leftarrow\min S$
18               return $w.\text{route}(v)$
19
Algorithm 1 route($v$)
Join.
The join algorithm (Algorithm 2) inserts the node inside the Continuous SLN based on its anchor local variable. The potential neighbor changes of the other nodes propagate through the network with the find_neighbors algorithm (Algorithm 3).
1
2 // code running at node $u$
3 if $\text{anchor}<u$ then
4   $S\leftarrow\{x\in N_{\text{anchor}_{u}}|\text{anchor}<x\}$
5
6   else
7     $S\leftarrow\{x\in N_{\text{anchor}_{u}}|x<\text{anchor}\}$
8
9     if $S=\emptyset$ then
10       $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
11       Update $D_{u}$
12       else
13         if $\text{anchor}<u$ then
14           $\text{succ}\leftarrow\min S$
15
16           else
17             $\text{succ}\leftarrow\max S$
18
19             $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
20             $N_{u}\leftarrow N_{u}\cup\text{succ}.\text{find\_neighbors}(u,h_{v},-\infty)$
21             Update $D_{u}$
22
Algorithm 2 join()
1
2 // code running at node $u$
3 if current_height = -$\infty$ then
4   $\text{anchor}_{u}\leftarrow v$
5
6  $N\leftarrow\emptyset$
7   if current_height $h_{u}$ then
8     $N\leftarrow\{u\}$
9     $N_{u}\leftarrow N_{u}\cup\{v\}$
10     Update $D_{u}$
11     if $u<v$ then
12       $N\leftarrow\{x\in N_{u}|x\leq v\text{ or }(h_{v}<D_{u}[x]\text{ and }h_{v}<h_{u})\}$
13       for $x\in N_{u}$ do
14         if $x<u\text{ and }h_{u}\leq D_{u}[x]\leq h_{v}$ then
15           $N\leftarrow N\cup x.\text{find\_neighbors}(v,h_{v},h_{u})$
16
17
18
19          else
20             $N\leftarrow\{x\in N_{u}|v\leq x\text{ or }(h_{v}<D_{u}[x]\text{ and }h_{v}<h_{u})\}$
21             for $x\in N_{u}$ do
22               if $u<x\text{ and }h_{u}\leq D_{u}[x]\leq h_{v}$ then
23                 $N\leftarrow N\cup x.\text{find\_neighbors}(v,h_{v},h_{u})$
24
25
26                 return NN
Algorithm 3 find_neighbors($v$, $h_{v}$, current_height)
Leave.
The leave algorithm (Algorithm 4, in this paper’s appendix) shows how an already-inserted node can leave the network. The potential neighbor changes propagate through the network with the delete_neighbors algorithm (Algorithm 5, in this paper’s appendix).
5 The Uniform SLN Protocol
After discussing the case of static SLN in the previous sections, we now model randomized P2P protocols. This section introduces a randomized, Continuous SLN called Uniform where each node chooses its height by picking a random real number in $(0,1)$ uniformly at random.
Definition 5.1 (Randomized Continuous SLN).
A randomized Continuous SLN is a probabilistic distribution over Continuous SLN.
Protocol 1 (Uniform).
Uniform is a randomized, continuous SLN that sets its heights to a real number in $(0,1)$ chosen uniformly at random.
The rest of this section presents the properties of Uniform with respect to routing path lengths and maximum degrees. We show that a Continuous SLN constructed with Uniform has an expected constant degree and an expected routing path length of $O(\log n)$. Also, we can rule out rare bad outcomes in practice: with high probability, the routing path length is $O(\log n)$ and the maximum degree is $O(\log n)$. Those results directly imply that the three primitive described in the previous section have a logarithmic message complexity with high probability. Interestingly, those results are quite general and hold for a vast family of continuous probability laws, namely, as soon as the probability that two heights are equal is zero.
Theorem 5.2 (Expected Path Length).
The routing path between any two nodes has length smaller than $2\ln n$ on expectation.
Proof 5.3.
Let $u$ and $v$ be two nodes. We note $X$ the random variable equal to the routing path length between $u$ and $v$. We note $X_{\text{up}}$ the number of height increases along the path, likewise $X_{\text{down}}$ is the number of path decreases. It holds that $X=X_{\text{up}}+X_{\text{down}}$. We will give an upper bound of $X_{\text{up}}$, the same upper bound will hold for $X_{\text{down}}$ by symmetry. Let $w$ be a node between $u$ and $v$ in the SLN, we note $1_{w}$ the indicator random variable equal to $1$ if (1) $w$ is part of the path between $u$ and $v$, and (2) if the height increases when hoping to $w$. It holds $X_{\text{up}}=\sum_{w\in[u+1,v]}1_{w}$.
We now compute $P[1_{w}=1]$ for any $\in[u+1,v]. ww$ is part of the increasing path if and only if the height of $w$ is strictly greater than every other heights in [u,w1][u,w-1]. The probability that two uniform random variables are equal is zero, hence there exists exactly one node with maximum height in $[u,w]$. The heights are independent and uniformly distributed, hence $w$ is the node with maximum height with probability 1/(|wu|+1)1/(|w-u|+1):
$\forall w\in[u+1,v],\quad P[1_{w}=1]=\frac{1}{|w-u|+1}$
Noticing that $\mathbb{E}[1_{w}]=P[1_{w}=1]$ and gathering the two previous equalities gives:
$\mathbb{E}[X_{\text{up}}]$ $=\sum_{w\in[u+1,v]}\frac{1}{|w-u|+1}=\sum_{i=2}^{|v-u|+1}\frac{1}{i}\leq\ln n$
The last inequality holds as the harmonic sum until $n$, $1+1/2+\dots+1/n$, is lower than $1+\ln n$. The claimed upper bound on the expected path length follows as the same reasoning can be done for $X_{\text{down}}$.
Theorem 5.4 (Routing Path Length with High Probability).
The routing path between any two nodes has length smaller than $2e\cdot\ln n$ with probability greater than $1-\frac{2}{n}$.
Proof 5.5.
Let $u$ and $v$ be two nodes such that $u<v$, and let $X_{\text{up}}$ be the random variable equal to the increasing routing path length between $u$ and $v$ . Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to $1$ if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 5.2).
We first show that all $1_{w}$ are independent random variables. Let $w$ and $w^{\prime}$ be two nodes such that $u<w<w^{\prime}$, it holds:
$P[1_{w^{\prime}}=1\mid 1_{w}=1]$ $=P[h_{x}<h_{w^{\prime}}\forall x\in[w+1,w^{\prime}-1]\cap\max_{y\in[u,w]}h_{y}<h_{w}^{\prime}]$
$=P[h_{y}<h_{w}^{\prime}\forall y\in[u,w^{\prime}-1]]$
$=P[1_{w^{\prime}}=1]$
Now that we showed independence, we use the following Chernoff’s bound: For $Y$ the sum of independent Bernoulli random variables (a Poisson trial) and $\mu=\mathbb{E}[Y]$, $\forall\delta>0$ it holds that:
$P(Y\geq(1+\delta)\mu)<\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{\mu}$
Replacing $Y=X_{\text{up}}=\sum_{w\in[u+1,v]}1_{w}$, $\mu=\ln n$ and $\delta=e-1$ gives:
$P(X_{\text{up}}\geq e\cdot\ln n)$ $<\frac{e^{\delta\cdot\ln n}}{(1+\delta)^{(1+\delta)\cdot\ln n}}$
$=\frac{n^{\delta}}{n^{(1+\delta)\cdot\ln(1+\delta)}}=\frac{1}{n}$
By symmetry, the same result holds for $X_{\text{down}}$. Finally, we obtain the desired result with a union bound:
$P(X\geq 2e\cdot\ln n)$ $\leq P(X_{\text{up}}\geq e\cdot\ln n\cup X_{\text{down}}\geq e\cdot\ln n)$
$\leq P(X_{\text{up}}\geq e\cdot\ln n)+P(X_{\text{down}}\geq e\cdot\ln n)$
$<\frac{2}{n}$
Theorem 5.6 (Expected Degree).
The expected degree of any node is lower than $4$.
Proof 5.7.
If $u$ and $w$ are two nodes, we define $N_{u}$ the random variable equal to the number of neighbors of $u$, and $1_{u,w}$ the indicator random variable equal to one if $u$ and $w$ are neighbors. It holds $N_{u}=\sum_{w\neq u}1_{u,w}$. Two nodes $u$ and $w$ with $u<w$ are neighbors if and only if $u$ and $w$ are the first and second highest nodes in $[u,w]$, hence it holds that $\mathbb{E}[1_{u,w}]=\frac{2}{|w-u|\cdot(|w-u|+1)}$. It therefore holds:
$\mathbb{E}[N_{u}]$ $=\sum_{w\in[1,n]\setminus\{u\}}\frac{2}{|w-u|\cdot(|w-u|+1)}$
$\leq 4\cdot\sum_{i=1}^{n/2}\frac{1}{i\cdot(i+1)}=4\cdot\sum_{i=1}^{n/2}\frac{1}{i}-\frac{1}{i+1}$
$=4\cdot(1-\frac{1}{n/2+1})\leq 4$
Theorem 5.8 (Maximum Degree with High Probability).
The maximum degree is lower than $4\ln n$ with probability greater than $1-\frac{1}{n^{3}}$.
Proof 5.9.
Let $u$, $w$ and $w^{\prime}$ be three different nodes. We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to $1$ if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively. We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 5.4.
We first establish that each node taken individually has a degree lower than $\ln n$ with probability greater than $1-\frac{1}{n^{4}}$. Let $N_{u}$ be the random variable equal to the number of neighbors of $u$, we use the Chernov bound for Poisson trials used in the proof of Theorem 5.4:
$P[N_{u}\geq 4\ln n]$ $<\left(\frac{e^{(\ln n)-1}}{(\ln n)^{\ln n}}\right)^{4}=\frac{(n/e)^{4}}{n^{4\ln\ln n}}<\frac{1}{n^{4}}$
This last equality holds when $n$ is greater than a given constant, which is not a problem as we are interested in cases when $n$ is large. We conclude with a union bound: the probability that at least one node has a degree greater than $4\ln n$ is upper bounded by $\sum_{v\in V}P[N_{v}\geq 4\ln n]\leq\frac{1}{n^{3}}$ and the claim holds.
6 Learning-Augmented Continuous SLN
In this section, we present LASLiN, a learning-augmented Continuous SLN protocol where each node can obtain a prediction of the future demand matrix at any time.
We first show that, for general demands, any randomized, demand-oblivious P2P protocol can be off by a factor of $O(\log n/\log\log n)$ compared to an optimum static SLN. This result provides a comparison point around which the consistency and robustness metrics of demand-aware algorithms can be evaluated.
Theorem 6.1.
We consider any randomized P2P protocol (Definition 5.1) that is ideal (Definition 2.4) and not learning-augmented (Definition 2.7). Then, there exists a demand matrix $W$ such that the considered protocol has a cost of $\Omega(\log n/\log\log n)\cdot|W|$ in expectation while an optimal static SLN protocol pays $|W|$, where $|W|$ is the total sum of the elements of $W$.
Proof 6.2.
As the considered protocol is ideal, there exists a constant $c$ such that the generated network has a maximum degree lower than $O(\ln^{c}n)$ with probability greater than 1O(1/n)1-O(1/n). Let $u$ be a node. We first cover the case when the maximum degree is lower than $\ln^{c}n$, all the considered probabilities and expected values are therefore conditional. In this case, we consider the disk centered around $u$ with smallest radius such that it contains no less than half of the nodes. The maximum distance between $u$ and a node in that disk is no lower than $O(\log n/\log\log n)$ as the network has a polylogarithmic maximum degree — this result can be easily obtained by upper bounding all the degrees by $O(\log^{c}n)$. There exists another node $v$ such that $v$ is at the border or outside of the disc with probability at least $1/2$. We consider the demand matrix that is zero everywhere except at the coefficient with row $u$ and column $v$. An optimal SLN protocol can clearly pay a cost of $1$, the considered randomized SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
Finally, we cover the cases where the maximum degree is higher than $\ln^{c}n$: That case only contributes with an additive constant to the distance between $u$ and $v$ as a distance is lower than $n$. We considered all the cases and the claim follows.
Protocol 2 (LASLiN).
LASLiN is a randomized, continuous SLN that sets its heights by adding the input, integer prediction in $1,\lceil\ln n\rceil]$, and a real number in $(0,1)$ chosen uniformly at random.
Theorem 6.3.
LASLiN is a randomized, Continuous SLN that is $O(1)$-consistent and $O(\log^{2}n)$-robust.
Proof 6.4.
We start by proving that the maximum degree in LASLiN is lower than $4\ln^{2}n$ with high probability. Let $u$ be a node, let $p_{u}\in[1,\lceil\ln n\rceil]$ be the predicted height of $u$ and let $r_{u}\in(0,1)$ be the uniform random noise added on top of $p_{u}$. We partition the neighbors of $u$ into $p_{u}$ subsets called $S(1),S(2)\dots S(p_{u}+1)$: For all $i\in[1,p_{u}]$, $S(i)$ contains the neighbors with height between $i$ and $i+1$, the last set $S(p_{u}+1)$ contains the neighbors with height greater than $p_{u}+1$. According to Theorem 5.8, each set from $S(1)$ to $S(p_{u})$ contains fewer than $4\ln n$ nodes with probability greater than $1-\frac{1}{n^{4}}$ (this is the result before the last paragraph in the proof) — the set $S(p_{u}+1)$ contains at most two nodes. We conclude with a union bound that $u$ has fewer than $4\ln^{2}n$ with probability greater than $1-\frac{1}{n^{3}}$, and that the maximum degree is lower than $4\ln^{2}n$ with probability greater than $1-\frac{1}{n^{2}}$ also with a union bound.
The $O(1)$-consistency is direct as the comparative optimal static SLN has its heights within a $O(\log n)$ factor without loss of generality as its maximum degree is bounded by $O(\log n)$ by assumption. We now prove the robustness claim. A routing path between any two nodes $u$ and $v$ has two phases: first ascending (the heights always increase) and then descending (the heights always decrease). For both phase, the number of different encountered integer heights is upper-bounded by $O(\log n)$. For each encountered integer height, the expected routing path length is upper-bounded by $O(\log n)$ (Theorem 5.2). We finish the proof with an upper bound of each distance in the total weighted path lengths: $\sum_{u,v\in V}W(u,v)\cdot d(u,v)\leq O(\log^{2}n)\cdot|W|$ and the optimal static SLN pays at least $|W|$.
We now define the notion of super-sparse matrix, which indicates a concentration of the weights of the matrix on a very small (logarithmic) number of coefficients.
Definition 6.5 (super-sparse matrix).
A matrix $W\in M_{n}(\mathbb{R})$ is super-sparse if there exists a constant $c\geq 1$ and a restricted subset of node pairs $S\subseteq V^{2}$ such that:
$|S|$ $\leq O(\log^{c}n)\qquad\text{ and }$
$\sum_{(u,v)\in S}W(u,v)$ $\geq\left(1-O(1/\log n)\right)\cdot|W|$
Theorem 6.6.
Assume that the demand matrix $W$ is super-sparse. We consider the protocol Binary LASLiN (B-LASLiN) where the input prediction $p$ can only have two values, $1$ or $2$. B-LASLiN is $O(\log\log n)$-consistent and $O(\log n)$-robust.
In asymptotics, the performances are therefore always better than any demand-oblivious protocol, with an exponential improvement in case of perfect predictions.
Proof 6.7.
We start by proving that the maximum degree of each node is lower than $O(\log n)$ with high probability. As seen in the proof of Theorem 6.3, each node $u$ has fewer than $p_{u}\cdot\ln n$ neighbors with high probability, the claim follows since $p_{u}\leq 2$.
We prove that the expected total weighted path length is lower than $O(\log n)\cdot|W|$ regardless of the input predictions. The proof is similar to the proof of Theorem 6.3 except, as the predictions are upper-bounded by a constant, we avoid the additional $\log n$ factor.
Finally, we prove that the expected total weighted path length is lower than $O(\log\log n)\cdot|W|$. We consider the following set of predictions: let $S$ be the node pairs from Definition 6.5, we give prediction $2$ to the nodes that have at least one coefficient from $S$ in their row or column in $W$, otherwise $1$. We call $V_{2}$ ($V_{1}$) the set of nodes that have prediction 22 (11). As $W$ is super sparse, it holds that $|V_{2}|\leq O(\log^{c}n)$.
$\sum_{u,v}W(u,v)\cdot d(u,v)$ $=\sum_{(u,v)\in S}W(u,v)\cdot d(u,v)+\sum_{(u,v)\in V^{2}\setminus S}W(u,v)\cdot d(u,v)$
$\leq\sum_{u,v\in V_{2}}W(u,v)\cdot d(u,v)+\sum_{(u,v)\in V^{2}\setminus S}W(u,v)\cdot d(u,v)$
$\leq O(\log\log n)\cdot|W|+O\left(\frac{1}{\log n}\right)\cdot|W|\cdot\ln n$
$=O(\log\log n)\cdot|W|$
To go from line 3 to 4 we used (1) Theorem 5.2 on the routing paths between the nodes of $V_{2}$ — there are $O(polylog(n))$ of those nodes, they have the largest heights of the SLN, so a path has expected length $\ln O(polylog(n))=O(\log\log n)$ — and on the routing paths from nodes in $V^{2}\setminus S$, and (2) we used Definition 6.5 to upper-bound the sum of coefficients in $W$ outside of $S$.
Figure 6: Comparison of the costs of multiple P2P networks against different demand matrix distributions. The demand matrices are generated based on the Zipfian distribution with varying skewness parameter $\zeta\in\{1.1,1.2\dots 2.0\}$, each coefficient is drawn independently and identically distributed. For each demand matrix distribution, we compare the cost of three P2P networks: the optimal static SLN, LASLiN, and the demand-oblivious Uniform SLN.
7 Experimental Evaluation
Demand matrices
In our experiments, we generate communication demand based on the Zipfian distribution [1], which receives a parameter $\zeta>0$, that indicates the distribution’s skewness. Roughly speaking, lower values of $\zeta$ result in a sparser demand matrix. The number of nodes in all experiments was set to $n=32$, and each experiment was repeated $30$ times.
Noise generation
As common in the literature [20], we study the robustness of our algorithms by introducing noise in the predicted data. The most used type of noise is additive, often gaussian [26] or lognormal [20]. This type of noise however almost did not impact the performance of our algorithms. Instead, we test two other models: the demand error recomputes a portion (x-axis) of the coefficients with the same Zipfian distribution, and the unit heights enforce that each node can receive a predicted height of $1$ with a certain probability (x-axis).
Results
In Figure 6 we evaluate the consistency of LASLiN in scenarios where the demand follows the Zipfian distribution, with varying parameter $\zeta\in[1.1,2]$, and where the prediction advice is perfect (i.e., there is zero noise in the demand matrix). We compare the total communication cost (as defined in (2.5)) of LASLiN to that of the static optimum SLN and the demand-oblivious SLN (Uniform) baselines. We can see that, in contrast to the demand-oblivious SLN, LASLiN obtains close-to-optimum total communication cost in all workloads, regardless of their spatial locality, i.e., LASLiN is consistent.
In Figure 7 we evaluate the robustness of LASLiN in scenarios with increasing demand noise (x-axis). We compare the total cost of LASLiN to that of the static optimum SLN, the demand-oblivious (Uniform) SLN, and the SLN baselines. We can observe that the SLN performs increasingly worse (has higher cost) the greater the amount of change (noise) in the input demand matrix. LASLiN, on the other hand, achieves consistently lower communication cost (than the former), regardless of the type of noise and regardless of the prediction fault probability, i.e., LASLiN is robust relative to both type and magnitude of prediction errors.
Figure 7: Cost of different P2P networks under varying noise amplitudes on the demand matrix predictions. Simulation results: robustness ($\zeta=1.2$)
8 Related Work
Improved designs of P2P networks. P2P networks have been the focus of computer scientists for a long while (see [19] for a detailed survey). In terms of P2P networks with constant degree, the notable contribution of [23] that extends the work of [13], giving a general P2P construction such that for a network with degree $d$ and $n$ nodes, you can have guaranteed path length of $\Theta(\log_{d}n)$. Lately, improving the design of networks given prior information about communication demand has grabbed the attention of researchers. In particular, the work of [7] set the scene for demand-aware network design, which was then explored more with particular focus on datacenter networking [11, 24, 12].
Works closest to ours are by [6, 10] in which the authors discuss a self-adjusting variant of skip list network: one that adjusts over time based on the incoming demand, which is in contrast with our work that focuses on an optimal skip list network with static height given the communication demand as the input. This is in particular important, as self-adjustments can introduce additional cost (e.g., in terms of delay) to the system. Also, the work of [21] discussed how a skip-list construction can be used for peer-to-peer communications, however they do not provide any theoretical upper bound for the communication time. Lastly, we point out to other extensions of skip list, namely Skip Graph [3] and SkipNet [16] that, to facilitate P2P communications, consider expected degree to be $O(\log n)$, in contrast to our design, which has a constant degree in expectation.
Learning-augmented algorithms. Given the rise of machine learning algorithms (that are prone to inherent unbounded failures), and after the pioneering work of [26] on ski rental & job scheduling, the focus has been shifted on studying consistency and robustness of algorithms augmented with the learned advice. Such an approach has been used in various applications so far222see a full list of recent works in the following website: algorithms-with-predictions.github.io, for example caching [20], online TSP [15], list update [8].
Some of the more recent works focus on improving single-source search data structures. Initially, the works of [18, 9] provided learning-augmented search trees. The works that are closest to our work are by [14, 30], which studied learning-augmented single-source skip lists. In contrast, we consider a more challenging model, where the focus is on pairwise communications rather than communication from a single source. Recently, we became aware of a not peer-reviewed manuscript [2] that aims to augment P2P networks with learned advice; however, this paper looks into improving self-stabilization and recovery of P2P networks, not improving communication delay.
9 Conclusion
In this work, introduced a continuous variant of Skip List Network (SLN), Continuous SLN. This variant is arguably easier to manipulate and analyze. In addition we show that we can use it to reach strong robustness properties even in an adversarial, fully distributed setting. Lastly, through experimental evaluation, we observed that our suggested algorithms outperform the theoretical results. As a future work, one can think of an alternative algorithm for OSSLN with a lower computational complexity.

### v2

@@
-LASLiN: A Learning-Augmented Peer-to-Peer Network
+A Learning-Augmented Overlay Network
 Abstract
-We introduce a learning-augmented peer-to-peer (P2P) network design that leverages the predictions of traffic patterns to optimize the network’s topology.
-While keeping formal guarantees on the standard P2P metrics (routing path length, maximum degree), we optimize the network in a demand-aware manner and minimize the path lengths weighted by the peer-to-peer communication demands.
-Our protocol is learning-augmented, meaning that each node receives an individual, possibly inaccurate prediction about the future traffic patterns, with the goal of improving the network’s performances.
-In this paper, we strike a trade-off between significantly improved performances when the predictions are correct (consistency) and polylogarithmic performances when the predictions are arbitrary (robustness).
-Our protocol builds upon a generalization of skip lists to networks.
-We have two main contributions.
-First, we consider the centralized setting and show that the problem of constructing an optimum static skip list network (SLN) is solvable in polynomial time and can be computed via dynamic programming.
-This problem is the natural demand-aware extension of the optimal skip list problem; in particular, we use it to generate the predictions of our learning-augmented protocol.
-Second, we introduce the Uniform P2P protocol which generalizes skip list networks (SLN) by allowing the nodes to have any real-valued height instead of discrete ones.
-We show that Uniform achieves state-of-the-art performances: logarithmic routing and maximum degree, both with high probability.
-We then use Uniform to build a learning-augmented P2P protocol in order to incorporate demand-awareness, leading to our main contribution, LASLiN.
-We prove that the performances of LASLiN are consistent with those of an optimum static SLN with correct predictions, and are at most a logarithmic factor off the state-of-the-art P2P protocols if the predictions are arbitrary wrong.
-For the special case of highly sparse demands, we show that LASLiN achieves an exponential improvement in case of correct predictions, and asymptotically matches the performances of state-of-theart, demand-oblivious P2P protocols in case of adversarial predictions.
-keywords:
-learning-augmented algorithms, peer-to-peer networks, skip list graphs
+This paper studies the integration of machine-learned advice in overlay networks in order to adapt their topology to the incoming demand.
+Such demand-aware systems have recently received much attention, for example in the context of data structures (Fu et al.
+in ICLR 2025, Zeynali et al.
+in ICML 2024).
+We in this paper extend this vision to overlay networks where requests are not to individual keys in a data structure but occur between communication pairs, and where algorithms have to be distributed.
+In this setting, we present an algorithm that adapts the topology (and the routing paths) of the overlay network to minimize the hop distance travelled by bit, that is, distance times demand.
+In a distributed manner, each node receives an (untrusted) prediction of the future demand to help him choose its set of neighbors and its forwarding table.
+This paper focuses on optimizing the well-known skip list networks (SLNs) for their simplicity.
+We start by introducing continuous skip list networks (C-SLNs) which are a generalization of SLNs specifically designed to tolerate predictive errors.
+We then present our learning-augmented algorithm, called LASLiN, and prove that its performance is (i) similar to the best possible SLN in case of good predictions ($O(1)$-consistency) and (ii) at most a logarithmic factor away from a standard overlay network in case of arbitrarily wrong predictions ($O(\log^{2}n)$-robustness, where $n$ is the number of nodes in the network).
+Finally, we demonstrate the resilience of LASLiN against predictive errors (ie, its smoothness) using various error types on both synthetic and real demands.
 1 Introduction
-Peer-to-peer (P2P) protocols are one of the key technological advancements that allowed decentralized and distributed systems to scale.
-A vast body of work has been devoted to optimize the efficiency of these protocols [19, 27, 3, 16, 17].
-Traditionally, the efficiency of a P2P protocol is estimated with two key metrics which are the routing path length between a pair of nodes and the maximum degree of each node in the network.
-State-of-the-art P2P protocols achieve both a logarithmic routing path length and a logarithmic maximum degree [29, 13].
-Up until recently, most work on P2P protocols focused on demand-oblivious worst-case scenarios, where no prior information about the possible future demand is known.
-However, a new line of work studies demand-aware networks [28, 5, 12] which can adapt their connectivity properties to better fit the current traffic.
-Noting that most networking traffics are sparse and partly predictable [4], the idea is to detect the few highly active nodes and move them topologically closer to reduce the overall delays.
-The model includes a demand matrix which indicates the amount of traffic between any two nodes in the network; A goal of a demand-aware protocol is then to minimize the total weighted path length [28].
-This paper presents a P2P network design where each node can adapt its set of neighbours to better fit the future traffic demands.
-We do not assume that the demand matrix is known by the nodes; Rather, we stick with the distributed nature of P2P protocols and assume that each node has access to a local prediction about the future demands — the prediction may only consist of partial information about the future traffic, an integer number in the case of this work.
-In our model, a node may obtain a new prediction at any time to use at its own discretion, independently from the other nodes.
-Crucially, the predictions are untrusted and we evaluate our P2P protocol in the learning-augmented setting with two metrics: consistency and robustness [20].
-The consistency of the algorithm is its performance when the predictions are always perfect, its robustness is its performance when the predictions are worst possible.
+Overlay networks are one of the key technological advancements that allowed decentralized and distributed systems to scale.
+A vast body of work has been devoted to optimize the efficiency of these networks Lua et al.
+(2005); Ratnasamy et al.
+(2001); Aspnes and Shah (2003); Harvey et al.
+(2003); Kiffer et al.
+(2021).
+Traditionally, the efficiency of an overlay network is estimated with two key metrics which are the routing path length between a pair of nodes and the maximum degree of each node in the network.
+State-of-the-art overlay networks achieve both a logarithmic routing path length and a logarithmic maximum degree Stoica et al.
+(2003); Fraigniaud and Gauron (2006).
+Up until recently, most work on overlay networks focused on demand-oblivious worst-case scenarios, where no prior information about the possible future demand is known.
+However, noting that most networking traffics are sparse and partly predictable Avin et al.
+(2020a), a new line of work studies demand-aware networks Schmid et al.
+(2016); Avin et al.
+(2020b); Figiel et al.
+(2025) which can adapt their connectivity properties to better fit the current traffic.
+On another front, there has been a recent surge in learning-augmented systems which, compared to traditional ones, take an additional parameter (the prediction) to improve their performances, with most recent examples in data structures Fu et al.
+(2025); Zeynali et al.
+(2024).
+There, the goal is to use a predicted demand for each node and adapt the system to better fit the incoming real demand.
+Since predicted and real demands may mismatch, they analyze what happens when the prediction is perfect (consistency), adversarial (robustness) and anything in between (smoothness, see Lykouris and Vassilvitskii (2018)).
+This work considers an overlay network with $n$ nodes and assume it faces a certain real demand.
+Like previous works on learning-augmented systems, our goal is to use a predicted demand to answer the overall demand with delays (cost) smaller than classic, prediction-oblivious systems.
+We use the framework of consistency and robustness to assess the performances of our overlay algorithm; for that we compare the performances of our learning-augmented algorithm with those of a Skip List Network (SLN, 1) whose heights are optimal with respect to the real demand.
+Unlike previous works however, our framework shows two crucial differences compared to other learning-augmented systems: our algorithm must (1) obey the standard constraints of overlay networks (Subsection 1.1), and (2) optimize on a pairwise demand representing the intensity of the exchanges between any two nodes (Subsection 1.1).
 1.1 Our Setup
-P2P Networks
-A P2P protocol is an algorithm distributed over multiple nodes that provides routing facilities between its member nodes.
-On top of a physical, static network, the P2P protocol keeps track of local routing information stored inside each node.
-This paper studies P2P network protocols in which each node $u$ holds a height variable $h_{u}$.
-In addition, we assume that each node holds a unique identifier (id) which defines a total order at any time: for any two different nodes $u$ and $v$, either $u<v$ or $v<u$.
-A Skip List Network (SLN[21] is a P2P design based on the well-known skip list data structure [25] which establishes the routing informations solely based on the heights and ids of the nodes.
-In a SLN, two nodes $u$ and $v$ form an edge, i.e., store each other’s id, if and only if there is no node $x$ such that $u<x<v$ and $h_{x}\geq\min(h_{u},h_{v})$ — see Figure 1.
-P2P Networks Constraints
-Following the overall consensus [13, 23] on P2P networks, we restrict our focus to P2P protocols with the following properties: (1) any node should be able to join or leave the network at any time, (2) the routing should be local, i.e., each new hop through the network is decided solely by one node, and (3) the routing should be oblivious, i.e., the id of the destination node is sufficient to route the message.
-Stated the above requirements, we want the network to still work at scale when the number of nodes in the network becomes large.
-Noting $n$ the number of nodes at the considered time, we aim at networks with:
-  • small routing path lengths. Given any two nodes $u$ and $v$, at all time, the routing path length between $u$ and $v$ is ideally logarithmic in $n$ with high probability.
-  • small degrees. At all time, the number of known nodes for any node, i.e., the maximum degree, is ideally logarithmic in $n$ with high probability.
-All the P2P protocols described of this paper respect those two criteria.
-To evaluate the performances of our P2P protocols, we compare them to idealized P2P protocols with the restriction that they need to have polylogarithmic path length and maximum degree (see Definition 2.4).
-Traffic Demand
-At a given time, we assume that the traffic demand between the nodes is modelled with a demand matrix $W\in M_{n}(\mathbb{R})$, where $W(u,v)$ is the frequency of communications from $u$ to $v$ [28].
-Learning-Augmented Protocols
-This paper introduces and studies learning-augmented P2P protocols.
-In the general case, a learning-augmented algorithm is an algorithm that takes an additional input, the prediction, to improve its performances.
-Crucially, the prediction is untrusted.
-This means that, with regard to the other inputs of the algorithm, the prediction may be perfect (all things being equal, no other prediction results in a better performance), adversarial, or anything in between.
-In our design of learning-augmented P2P protocols, we assume that each node can obtain a prediction at any time, independently from one another.
-The origin of the prediction is not the focus of this paper; it may come from a centralized server handing out traffic predictions to the nodes, or it may be locally computed by each node based on its own traffic history.
-In this work, the input prediction is an integer interpreted as the height of the node in the SLN.
-1.2 Contributions and Technical Novelty
+Network and Demand.
+We model an overlay network by an undirected graph with $n$ nodes.
+An edge of the graph means that the two end-nodes can communicate directly without the help of other nodes, ie, they hold each other’s ids.
+This graph is an abstract representation of the routing possibilities between the nodes and may be different from the underlying, physical network that carries the communications.
+At the considered time, we model a demand between the nodes with a demand matrix $W\in M_{n\times n}(\mathbb{R})$, where $W(u,v)$ is the amount of communications from $u$ to $v$.
+Overlay Network Algorithms.
+An overlay network algorithm is an algorithm distributed over the nodes of the network that defines the connectivity between overlay nodes and how routing can be performed.
+On top of the physical, static network, the overlay algorithm keeps track of local routing information stored inside each node.
+Prediction Model.
+The initialization procedure of our overlay algorithm (run right before the join procedure) takes a prediction as an additional argument (see also Aradhya and Scheideler (2025)).
+As a newcomer node joins the overlay network, the prediction influences its neighborhood in the network to (ideally) fit the incoming demand best.
+In the case of this paper, the prediction is the (integer) height of the node in the considered skip list network.
+Adaptation to a changing demand (and thus changing prediction) is performed by leaving and re-joining the network.
+Consistency, Robustness, and Smoothness.
+The predictions are untrusted and should not be followed blindly by the nodes.
+In this context, we evaluate the performances of our algorithm with the two most frequently studied metrics called consistency and robustness Lykouris and Vassilvitskii (2018) (7).
+In more details, we derive overlay algorithms that provably improve their performance in case of globally accurate predictions (consistency), but still present good performances (poly-logarithmic in $n$ with high probability) in case of arbitrary predictions, even adversarial ones (robustness).
+In our distributed context however, we observe that the consistency metric may be too weak and its assumption that all nodes have good predictions unrealistic.
+We therefore consider a third metric, smoothness, which indicates the performances along the spectrum from consistency to robustness; in particular, what happens when the prediction is good but with a slight noise.
+We provide extensive analysis of the smoothness of our algorithm tested on synthetic and real demands and with different types of errors (Section 6).
+1.2 Contributions
 The contributions of this paper are threefold.
-First, we formulate the Optimum Static Skip List Network (OSSLN) problem and prove it can be solved in polynomial time, with an algorithm based on dynamic programming (Theorem 3.3).
-This result extends the previous work of [22] to the case with pairwise demands in line with recent demand-aware results [28].
-Later on, this algorithm demonstrates that the task of finding optimal predictions for our learning-augmented P2P protocol is also polynomial.
-012312345678
-Figure 1: A representation of a Skip List Network (SLN) with eight nodes.
-The nodes have ids ranging from $1$ to $8$.
-The height of a rectangle equals the height of its corresponding node; Notice that all heights have integer values as this is an SLN.
-The double-arrowed, horizontal lines are the edges of the SLN.
-0112345678
-Figure 2: A continuous SLN with real-valued heights in $(0,1)$.
+First, we formulate the Optimum Static Skip List Network (OSSLN) problem and prove it can be solved in polynomial time, with an algorithm based on dynamic programming (Theorem 2).
+This result extends the previous work of Martinez and Roura (1995) to the case with pairwise demands in line with recent demand-aware results Schmid et al.
+(2016).
+This algorithm demonstrates that the task of finding optimal predictions for our learning-augmented overlay algorithm is polynomial, which allowed us to run our evaluations.
+As our second contribution, we present a new overlay design called Continuous Skip List Network (C-SLN, 9), which generalizes Skip List Networks Avin et al.
+(2020c) (SLN, 1) and the classic skip lists Pugh (1990).
+We show that a C-SLN offers a functioning overlay algorithm in that we implement the three basic primitives join, leave and route.
+We then present a specific (without prediction) C-SLN called Uniform that draws its heights uniformly at random in the interval $[0,1]$.
+We show that Uniform matches the performances of state-of-the-art overlay algorithms with both routing path lengths and maximum degrees in $O(\log n)$ with high probability.
+As our third contribution, we present a learning-augmented overlay algorithm, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform algorithm with a predicted SLN to strike consistency and robustness guarantees (7).
+For each node, LASLiN takes as prediction an integer between $1$ and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $[0,1]$ uniformly at random and adds it to the integer heights to obtain its final heights as a C-SLN; it may then join the network using the join() procedure.
+We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $O(\log^{2}n)$ factor away from the optimum static SLN in case of adversarial predictions ($O(\log^{2}n)$-robust).
 0123412345678
-Figure 3: A Continuous SLN obtained by adding node-by-node the heights of the SLN in Figure 1 and the heights of the Continuous SLN in Figure 2.
-This is a possible outcome of LASLiN where the predicted heights are those of Figure 1 and the uniform, random heights are those of Figure 2.
-01212345678
-Figure 4: A SLN with all equal heights, effectively turning the SLN into a simple chained list.
-This SLN has a linear routing path length.
-This worst-case performance happens when all the nodes get the same predicted height.
-01212345678
-Figure 5: The addition of continuous, random heights restores the good performances.
-As our second contribution, we present a new P2P design called Continuous Skip List Network (Continuous SLN, Definition 4.1), which generalizes Skip List Networks [6] (SLN, Definition 2.1) and skip lists [25].
-We show that a SLN with real-valued heights enables a functioning P2P protocol in that we implement the three basic primitives route, join and leave with state-of-the-art, logarithmic message complexity.
-We note that all known P2P designs that are based on skip lists [16, 14, 3] historically perceive it as chained lists stacked on each other and explicitly use that discrete nature in the protocol implementation, we show that this is not necessary.
-We then present a Continuous SLN protocol called Uniform that draws its heights uniformly at random in the interval $(0,1)$ and show that Uniform matches the performances of state-of-the-art P2P protocols with both routing path lengths and maximum degrees lower than $O(\log n)$ with high probability.
-As our third contribution, we present a learning-augmented P2P protocol, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform protocol with a predicted SLN to strike consistency and robustness guarantees (Definition 2.8).
-For each node, LASLiN takes as prediction an integer between $1$ and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $(0,1)$ uniformly at random and adds it to the integer heights to obtain its final heights as a Continuous SLN.
-We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $\ln n$ factor off a demand-oblivious P2P protocol with arbitrary predictions ($O(\log^{2}n)$-robust).
-Finally we show that, in the special case of a very sparse demand matrix (see Definition 6.5), the version of LASLiN with binary predictions always outperforms any demand-oblivious P2P protocol with high probability: LASLiN obtains an exponentially better cost with correct predictions while it asymptotically matches the cost of the state-of-the-art P2P protocols for arbitrary predictions (see Theorem 6.6).
-We provide simulations to assess the performance of LASLiN on Zipfian demand matrices.
+Figure 1: A SLN with eight nodes. The node ids range from $1$ to $8$. The double-arrowed, horizontal lines are the edges of the SLN.
+0123412345678
+Figure 2: A C-SLN obtained by adding each height of Figure 2 with a random number in $[0,1]$. This is a possible outcome of LASLiN with predicted heights in green.
 2 Model and Metrics
-In this section, we formally define our model and the metrics we use to evaluate our P2P protocols.
-Definition 2.1 (SLN).
+In this section, we formally define our model and the metrics we use to evaluate our overlay algorithms.
+Definition 1 (SLN).
 A Skip List Network (SLN) $\mathcal{N}$ is a triple $(V,h,E)$ where:
   • $V$ is a totally ordered set of node ids;
-  • $h$ is a height function $h:V\longrightarrow\mathbb{N}^{+}$;
-  • $E=\{\{u,v\}\;|\;\forall x\in V\;s.t.\;u<x<v,\;h_{x}<\min(h_{u},h_{v})\}$.
-Definition 2.2 (Neighbor and Degree).
-We say that two nodes $u$ and $v$ are neighbors in a SLN $\mathcal{N}=(V,h,E)$ if $\{u,v\}\in E$. The degree of a node $u$ is its number of neighbors.
-We redefine the greedy, local routing path proper to SLN.
-Definition 2.3 (Routing Path).
-Let $\mathcal{N}=(V,h,E)$ be a SLN and $u,v\in V$ be two different nodes.
-The routing path from $u$ to $v$ in $\mathcal{N}$ is the sequence of edges of $E$ $s=e_{1},e_{2}\dots e_{k}$ such that: (1) $u\in e_{1}$, (2) $v\in e_{k}$, and (3) $\forall i\in[k]$, if $u<v$ then $e_{i}=\{x,y\}$ where $y$ is the largest node neighbor of $x$ such that $x<y\leq v$, otherwise if $v<u$ then $e_{i}=\{x,y\}$ where $y$ is the smallest node neighbor of $x$ such that $v\leq y<x$.
-More intuitively, the routing path goes to the next neighbor that is the closest to the destination (in terms of ids) without “flying over” the destination.
-Definition 2.4 (Ideal P2P Protocol).
-We say that a (potentially randomized) P2P protocol is Ideal if, at all times, the routing path length between two nodes and the maximum degree are polylogarithmic on $n$ with high probability, where $n$ is the number of nodes at the considered time.
-A quantity is polylogarithmic when there exists positive constants $c$ and $d$ such that the quantity is smaller than $d\cdot\log^{c}n$.
-Definition 2.5 (Cost).
+  • $h$ is a height function $h:V\rightarrow\mathbb{N}^{+}$;
+  • $E$ is the set of edges such that $\{u,v\}\in E$ if and only if
+    $\forall x\in V\;\text{ s.t. }\;u<x<v,\quad h_{x}<\min(h_{u},h_{v})$
+We redefine hereafter the greedy, local routing path specific to SLN.
+Intuitively, the routing path goes to the next neighbor that is the closest to the destination (in terms of ids) without “flying over” the destination.
+Definition 2 (Routing path).
+[…]
+Definition 3 (Cost).
 Let $\mathcal{N}=(V,h,E)$ be an SLN and $W$ be a demand matrix. The cost function is:
@@
 where $d_{\mathcal{N}}(u,v)$ is the routing path length between $u$ and $v$ in $\mathcal{N}$.
-Definition 2.6 (Optimum Static Skip List Network (OSSLN) problem).
-Given a set of node ids $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\longrightarrow\mathbb{N}^{+}$ that minimizes the cost of the SLN $\mathcal{N}=(V,h,E)$.
-Definition 2.7 (Learning-augmented P2P protocol).
-A learning-augmented P2P protocol ALG is a P2P protocol whose initialization routine takes an additional parameter $p$, the prediction.
-In our model, a node that wishes to update its prediction as it already joined the network simply leaves, re-intializes and joins again.
-Definition 2.8 (Consistency and Robustness).
-We say that a learning-augmented SLN protocol is
-  • $\alpha$-consistent if, for all demand matrices, there exists an assignment of predictions to the nodes such that the expected cost of the protocol is at most $\alpha$ times the cost of an ideal optimum static SLN.
-  • $\beta$-robust if, for all demand matrices and for all assignment of predictions, the expected cost of the protocol is at most $\beta$ times the cost of an ideal optimum static SLN.
-3 Optimum Static Skip List Network (OSSLN)
-Given a set of node ids $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\longrightarrow\mathbb{N}^{+}$ that minimizes the total communication cost (as defined in (2.5)) of the SLN $\mathcal{N}=(V,h,E)$ (as defined in (2.1)):
-$OSSLN(V,W)=\min_{h}\;Cost(\mathcal{N},W).$ (1)
-In [28], the authors presented an algorithm based on dynamic programming to find the Optimum static distributed Binary Search Tree (OBST) in polynomial time ($O(n^{3})$).
-This algorithm leverages a crucial property: once a vertex $v$ is chosen as the root of a tree induced by the vertices in a given interval, the subproblems induced by the vertex id intervals to the left and to the right of $v$ can be solved independently.
-The OSSLN problem exhibits a similar structure, but with crucial difference.
-If we define a pivot vertex $x\in[l,r]$, $1\leq l\leq r\leq n,$ such that $h_{x}=\max_{w\in[l,r]}h_{w}$, $x$ also partitions the OSSLN problem into two subproblems (on the intervals $[l,v)$ and $(v,r]$).
-However, a critical distinction arises when $h_{x}<\max_{w\in[l,r]}h_{w}$, in which case $x$ will not be a part of all the paths between vertex pairs $(u,v)$ $u\in[l,x),v\in(x,r]$.
-Specifically, for a vertex $x\in[l+1,r-1]$ such that $h_{l},h_{r}>h_{x}$, any path between vertices $(u,v)$ where $u\leq l$ and $v\geq r$ does not include $x$.
+Definition 4 (Optimum Static Skip List Network (OSSLN) problem).
+Given a set of nodes $V$ and a demand matrix $W$, the objective of the Optimum Static SLN (OSSLN) problem is to find a height assignment $h:V\rightarrow\mathbb{N}^{+}$ that minimizes the cost of the SLN $\mathcal{N}=(V,h,E)$.
+Definition 5 (Overlay Algorithm).
+An overlay algorithm is a distributed algorithm running on each node of the network, which provides three procedures:
+  • initialization: Executed before a node joins the network. In this paper, sets the height of each node in the SLN.
+  • join: Executed to join the network. Takes as argument the id of a node that is already in the network.
+  • leave: Executed to leave the network.
+Definition 6 (Learning-Augmented Overlay Algorithm).
+A learning-augmented overlay algorithm ALG is an overlay algorithm whose initialization procedure takes an additional parameter, the input prediction.
+In this paper, the input prediction consists of an estimated height assignment for each node.
+Definition 7 (Consistency and Robustness).
+We say that a learning-augmented C-SLN algorithm is
+  • $\alpha$-consistent if, for all demand matrices $W$ and all set of nodes $V$, there exists a set of predictions (assignment of heights) such that the expected cost of the algorithm is at most $\alpha$ times the cost of an optimal optimum static SLN for $V$ and $W$.
+  • $\beta$-robust if, for any demand matrix $W$, any set of nodes $V$, and any set of prediction (assignment of heights), the expected cost of the algorithm is at most $\beta$ times the cost of an optimal optimum static SLN for $V$ and $W$.
+3 Optimum Static Skip List Network
+Given a set of nodes $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\rightarrow\mathbb{N}^{+}$ that minimizes the total communication cost (3) of the SLN $\mathcal{N}=(V,h,E)$ (1):
+$\textsc{OSSLN}(V,W)=\min_{h}\;Cost(\mathcal{N},W).$ (1)
+where $\mathcal{N}$ is an SLN with node set $V$ and heights $h$.
+In Schmid et al.
+(2016), the authors presented a dynamic programming algorithm to find the Optimum static distributed Binary Search Tree (OBST) in polynomial time ($O(n^{3})$).
+This algorithm leverages a crucial property: once a node $v$ is chosen as the root of a tree induced by the nodes in a given interval, the subproblems induced by the node id intervals to the left and to the right of $v$ can be solved independently.
+The OSSLN problem exhibits a similar structure, but with a crucial difference.
+If we define a pivot node $x\in[l,r]$, $1\leq l\leq r\leq n,$ such that $h_{x}=\max_{w\in[l,r]}h_{w}$, $x$ also partitions the OSSLN problem into two subproblems (on the intervals $[l,v)$ and $(v,r]$).
+However, a critical distinction arises when $h_{x}<\max_{w\in[l,r]}h_{w}$, in which case $x$ will not be a part of all the paths between node pairs $(u,v)$ $u\in[l,x),v\in(x,r]$.
+Specifically, for a node $x\in[l+1,r-1]$ such that $h_{l},h_{r}>h_{x}$, any path between nodes $(u,v)$ where $u\leq l$ and $v\geq r$ does not include $x$.
 This changes the optimal substructure of the problem, requiring a more sophisticated dynamic programming approach.
-Given a SLN $\mathcal{N}$, we define a path $P_{h}(l,r)$ to be the set of vertices in the routing path between the vertices $l$ and $r$ (as defined in (2.3)), excluding $l$, i.e., $l\notin P_{h}(l,r)$.
+Given a SLN $\mathcal{N}$, we define a path $P_{h}(l,r)$ to be the set of nodes in the routing path between the nodes $l$ and $r$ (2), excluding $l$, i.e., $l\notin P_{h}(l,r)$.
 This is a crucial simplification that allows the problem to be decomposed.
@@
 W.l.o.g., we assume that $P_{h}(l,r)=P_{h}(r,l)$.
-Next, we define $L(x,h)=\min\{i\leq x\;|\;h_{j}\leq h_{x},\;\forall j\in[i,x]\}$, i.e., the vertex of minimum id between the set of all (consecutive) vertices of height lower or equal to $h_{x}$, to the left of $x$, and, analogously, $R(x,h)=\max\{i\geq x\;|\;h_{j}\leq h_{x},\;\forall j\in[x,i]\}$, i.e., the vertex with maximum id between the set of all vertices of height lower or equal to $h_{x}$, to the right of $x$.
-A vertex $x\in V$ belongs to a path $P_{h}(u,v)$ if and only if there does not exist any pair of vertices $(u^{\prime},v^{\prime})$ such that $u\leq u^{\prime}<x$, $x<v^{\prime}\leq v$, and (both) $h_{u^{\prime}},h_{v^{\prime}}>h_{x}$.
+Next, we define $L(x,h)=\min\{i\leq x\;|\;h_{j}\leq h_{x},\;\forall j\in[i,x]\}$, i.e., the node of minimum id between the set of all (consecutive) nodes of height lower or equal to $h_{x}$, to the left of $x$, and, analogously, $R(x,h)=\max\{i\geq x\;|\;h_{j}\leq h_{x},\;\forall j\in[x,i]\}$, i.e., the node with maximum id between the set of all nodes of height lower or equal to $h_{x}$, to the right of $x$.
+A node $x\in V$ belongs to a path $P_{h}(u,v)$ if and only if there does not exist any pair of nodes $(u^{\prime},v^{\prime})$ such that $u\leq u^{\prime}<x$, $x<v^{\prime}\leq v$, and (both) $h_{u^{\prime}},h_{v^{\prime}}>h_{x}$.
 Finally, given an arbitrary $x\in V$, we define $P_{x}(h)$ as the set of all pairs $(u,v)\in V\times V$, such that $x\in P_{h}(u,v)$.
-Lemma 3.1.
+Lemma 1.
 If $h$ is a solution for the $OSSLN(V,W)$, then $\forall x\in V$ we have that:
-$P_{x}(h)$ $=$ $\{(u,v)\mid u<x\;\wedge\;v\in[x,R(x,h)]\}\;\cup$
-$\{(u,v)\mid u\in[L(x,h),x)]\;\wedge\;v>R(x,h)\}$
-Proof 3.2.
+$P_{x}(h)$ $=$ $\{(u,v)\mid(u<x)\wedge v\in[x,R(x,h)]\}\cup$
+$\{(u,v)\mid u\in[L(x,h),x)]\wedge(v>R(x,h))\}.$
+Proof.
 If $u<x$ and $x\leq v\leq R(x,h)$, then, by definition of $R(x,h)$, $h_{x}\geq h_{v}\geq\min(h_{u},h_{v})$.
-Since $x$ lies between $u$ and $v$ and there are no edges “above" $h_{x}$ in the interval $[u,v]$ (i.e., $\nexists(w,z)\in E$, s.t.
+Since $x$ lies between $u$ and $v$ and there are no edges “above” $h_{x}$ in the interval $[u,v]$ (i.e., $\not\exists(w,z)\in E$, s.t.
 $h_{w}>h_{x}$ and $h_{z}>h_{x}$, $w,z]\subseteq[u,v]$), $x$ must be a part of $P_{h}(u,v)$.
 Similarly, if $L(x,h)\leq u<x$ and $v>R(x,h)$, then, again by definition of $L(x,h)$, $h_{x}\geq h_{u}\geq\min(h_{u},h_{v})$.
-Since $x$ lies between $u$ and $v$ and there are no edges "above" $h_{x}$ in the interval $[u,v]$, $x$ must be a part of $P_{h}(u,v)$
+Since $x$ lies between $u$ and $v$ and there are no edges ”above” $h_{x}$ in the interval $[u,v]$, $x$ must be a part of $P_{h}(u,v)$
 It remains to show that no other pair $(u,v)$ belongs to $P_{x}(h)$.
@@
 $|P_{h}(u,u^{\prime})\cup\{v^{\prime}\}\cup P_{h}(v^{\prime},v)|<|P_{h}(u,x)\cup\{x\}\cup P_{h}(x,v)|.$
-Because $P_{h}(u,v)$ is the shortest path from $u$ to $v$, it follows that $x\notin P_{h}(u,v)$.
+Because $P_{h}(u,v)$ is the shortest path from $u$ to $v$, it follows that $x\notin P_{h}(u,v)$. ∎
 Since $d(u,v)=|P_{h}(u,v)|$, we can now rewrite the total cost of an SLN $\mathcal{N}=(V,h,E)$ as:
-$\text{Cost}(\mathcal{N},W)$ $=$ $\sum_{u,v\in V}W(u,v)\cdot d(u,v)$
-$=\sum_{u,v\in V}\sum_{x\in P_{h}(u,v)}W(u,v)$ $=$ $\sum_{x\in V}\sum_{(u,v)\in P_{x}(h)}W(u,v).$
-This means that the total cost is equal to the sum of contributions of all vertices $x\in V$, where the contribution of a vertex $x$ to the total cost is defined as:
+$Cost(\mathcal{N},W)$ $=$ $\sum_{u,v\in V}W(u,v)\cdot d(u,v)$
+$=$ $\sum_{u,v\in V}\sum_{x\in P_{h}(u,v)}W(u,v)$
+$=$ $\sum_{x\in V}\sum_{(u,v)\in P_{x}(h)}W(u,v).$
+This means that the total cost is equal to the sum of contributions of all nodes $x\in V$, where the contribution of a node $x$ to the total cost is defined as:
 $\sum_{(u,v)\in P_{x}(h)}W(u,v)$ (2)
-OSSLN with Bounded Height: In the OSSLN with bounded height, the height assignment is of the form $h:V\longrightarrow[1,h_{\max}]$, where $h_{\max}$ is a parameter, typically bounded by $O(\log{n})$ or $O(1)$.
-Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$.
-Let $nl$, $1\leq nl\leq l$, denote the leftmost vertex id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let $nr$, $r\leq nr\leq n$, denote the rightmost vertex id, such that $\forall v\in[r,nr],\,h_{v}\leq h$.
-Finally, let $\text{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all vertices $v\in[l,r]$.
-Theorem 3.3.
-If $x\in[l,r]$ is a candidate pivot vertex with height $h\in[1,h_{\max}]$, then the following recurrence relation holds $\forall(nl,l,x,r,nr)$, $1\leq nl\leq l\leq x\leq r\leq nr\leq n$:
-$\text{OSSLN}(l,r,nl,nr,h,W)=\min\Bigg{\{}\text{OSSLN}(l,r,l,r,h-1,W),$
-$\min_{x\in[l,r]}\Bigg{[}\text{OSSLN}(l,x-1,nl,nr,h,W)+$
-$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
-$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\Bigg{]}\Bigg{\}}.$
-Proof 3.4.
-The proof is by strong induction on the size of the interval $l,r]\subseteq[1,n]$:
-Note that, by definition, $nl=L(x,h)$ and $nr=R(x,h)$ for any possible $h$.
-Furthermore, by construction, for an interval induced by $l$, $r$, nlnl, nrnr with bounded height $h$, assigning a pivot vertex with height equal to h1h-1, we have $nl^{\prime}=l$ and $nr^{\prime}=r$.
-Base case $(l=r)$: In this case, we have that $(x=l)$ and the set of paths $x$ belongs to is given by Lemma 3.1.
-By setting $nl=L(x,h)$ and $nr=R(x,h)$ in the proof, the vertex contribution to the total cost is given by:
-$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)$ (4)
-Induction step: For $(h=1)$, we can disregard the first half of the recursion, where we assign the pivot vertex a height lower than $h$. The total contribution of all vertices to the cost becomes:
-$\sum_{x\in[l,r]}\left(\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\right),$
-which, from the base case, is precisely what the second half of the recurrence equation will return for any choice of the pivot vertex.
-Let $1\leq l<r\leq n$ and $h\in[1,h_{\max}]$.
-Assume that the recurrence equation is true for all $h^{\prime},l^{\prime},r^{\prime}$ such that $(h^{\prime}<h)\leavevmode\nobreak\ \wedge\leavevmode\nobreak\ (r^{\prime}-l^{\prime})<(r-l)$.
-From previous step also assume that the equation is true for $(r^{\prime}-l^{\prime})=(r-l)$ with height $(h^{\prime}<h)$.
-Since, when assigning the pivot vertex with height lower than $h$, by construction, we have that $(nl=l)$ and $(nr=r$).
-The minimum cost for the solution for any $nl$ and $nr$ is the minimum between keeping the height of the pivot, $(h_{x}=h)$, or decreasing it, $h_{x}\leq(h-1)$, which is given by the minimum between:
-$\min_{x\in[l,r]}\Bigg{[}\text{OSSLN}(l,x-1,nl,nr,h,W)+$
-$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
-$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\Bigg{]}$
-and OSSLN(l,r,l,r,h1,W)OSSLN(l,r,l,r,h-1,W), which is precisely what we want to prove.
-If we pre-compute an auxiliary matrix of aggregate weights ($W^{\prime}(x,y)=\sum_{z\leq y}W(x,z),\forall(x,y),1\leq x\leq y\leq n$) in time $O(n^{3})$, then the above dynamic programming algorithm has complexity $\mathcal{O}(n^{5}\cdot h_{\max})$.
-OSSLN with Unbounded height (OSSLNU): We now present a dynamic programming algorithm for the OSSLN problem with unbounded heights, i.e., when the height assignment is of the form $h:V\longrightarrow[1,n]$.
-We show that, in this particular case, the algorithm’s complexity can be reduced from $\mathcal{O}(n^{5}\cdot h_{\max})$ to $\mathcal{O}(n^{3})$.
-A crucial observation is that, for any solution to the OSSLNU problem, there exists a solution of cost less than or equal to that of the former, in which, $\forall[l,r],1\leq l\leq r\leq n$, there is a unique pivot vertex, i.e., a unique vertex whose height is the maximum height in that interval.
-Lemma 3.5.
-Let $h$ be an optimum solution to $OSSLNU(V,W)$.
-Then, there must exist a solution $h^{\prime}$ such that $Cost((V,h^{\prime},E^{\prime}),W)\leq Cost((V,h,E),W)$, where, for any interval $l,r],1\leq l\leq r\leq n,$ there is exactly one vertex $x\in[l,r]$ with height $h^{\prime}_{x}=\max_{w\in[l,r]}h^{\prime}_{w}$.
-Proof 3.6.
+OSSLN with Unbounded height (OSSLNU): We now present a dynamic programming algorithm for the OSSLN (4). Furthermore, we show that the algorithm’s complexity can be reduced to $\mathcal{O}(n^{3})$.
+A crucial observation is that, for any solution to the OSSLNU problem, there exists a solution of cost less than or equal to that of the former, in which, $\forall[l,r],1\leq l\leq r\leq n$, there is a unique pivot node, i.e., a unique node whose height is the maximum height in that interval.
+Lemma 2.
+Let $h$ be an optimum solution to $\text{OSSLNU}(V,W)$.
+Then, there must exist a solution $h^{\prime}$ such that $Cost((V,h^{\prime},E^{\prime}),W)\leq Cost((V,h,E),W)$, where, for any interval $l,r],1\leq l\leq r\leq n,$ there is exactly one node $x\in[l,r]$ with height $h^{\prime}_{x}=\max_{w\in[l,r]}h^{\prime}_{w}$.
+Proof.
 Suppose, for contradiction, that any possible optimal solution contains at least one interval $[l,r]$ with at least two vertices $x,y\in[l,r]$ with $h_{x}=h_{y}=\max(h_{l,r})$ and $x\neq y$.
@@
 Hence, the lemma holds.
-In OSSLNU, because we are only interested in the relative order of vertex heights, rather than in their absolute values, we have that $(nl=l)$ and $(nr=r)$, $\forall[l,r],1\leq l\leq r\leq n$, and the height parameter $h$ becomes unnecessary.
-This greatly simplifies the recurrence relation, reducing the complexity to $\mathcal{O}(n^{3})$.
-Theorem 3.7.
+∎
+In OSSLNU, because we are only interested in the relative order of node heights, rather than in their absolute values, we have that $(nl=l)$ and $(nr=r)$, $\forall[l,r],1\leq l\leq r\leq n$, and the height parameter $h$ becomes unnecessary.
+This simplifies the recurrence relation, reducing the complexity to $\mathcal{O}(n^{3})$.
+Theorem 1.
 Consider an interval $[l,r]$, $1\leq l\leq r\leq n$.
-Let $\text{OSSLNU}(l,r,W)$ denote the minimum total cost contributed by all vertices $v\in[l,r]$.
-Let $x\in[l,r]$ be a candidate pivot vertex.
+Let $\text{OSSLNU}(l,r,W)$ denote the minimum total cost contributed by all nodes $v\in[l,r]$.
+Let $x\in[l,r]$ be a candidate pivot node.
 Define $l$ as the leftmost index such that $\forall x,\in[l,r],\forall v\in[l,x],\,h_{v}\leq h_{x}$, and similarly define $r$ as the rightmost index such that $\forall v\in[x,r],\,h_{v}\leq h_{x}$.
-Then:
-$\text{OSSNU}(l,r,W)=\min_{x\in[l,r]}\Big{(}\text{OSSNU}(l,x-1,W)+$
-$\text{OSSNU}(x+1,r,W)+$
-$\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)\Big{)}.$ (5)
-Proof 3.8.
+Then, the recurrence relation $\text{OSSNU}(l,r,W)=$
+$\min_{x\in[l,r]}\Big(\text{OSSNU}(l,x-1,W)+\text{OSSNU}(x+1,r,W)+$
+$\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)\Big).$
+Proof.
 The proof is by induction on the size of the interval $[l,r]$.
-Note that, by Lemma 3.5, once a pivot vertex $x\in[l,r]$ is chosen as the highest vertex within the interval $[l,r]$, we can assume, without loss of generality, that all vertices in the optimal solutions to the left subinterval [l,x1][l,x-1] and to the right subinterval $[x+1,r]$ have heights strictly lower than $h_{x}$.
+Note that, by Lemma 2 once a pivot vertex $x\in[l,r]$ is chosen as the highest vertex within the interval $[l,r]$, we can assume, without loss of generality, that all vertices in the optimal solutions to the left subinterval [l,x1][l,x-1] and to the right subinterval $[x+1,r]$ have heights strictly lower than $h_{x}$.
 Therefore, for a solution $h$ to $\text{OSSLNU}(l,r,W)$, it follows that $l=L(x,h)$ and $r=R(x,h)$.
 Base Case $(l=r)$: In this case, the interval contains a single vertex $x=l=r$.
-In this case, the cost reduces to the contribution of $x$ to the solution cost, which, from Lemma 3.1, and $L(x,h)=l$ and $R(x,h)=r$ as we demonstrated above, it follows that this expression corresponds exactly to:
+In this case, the cost reduces to the contribution of $x$ to the solution cost, which, from Lemma 1, and $L(x,h)=l$ and $R(x,h)=r$ as we demonstrated above, it follows that this expression corresponds exactly to:
 $\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)$ (3)
@@
   • The optimal cost for the right subinterval $[x+1,r]$,
-  • The cost of all paths that include $x$, which by Lemma 3.1 is given by the Eq.(6).
-Since we minimize this total over all pivot choices $x\in[l,r]$, the recurrence holds.
-4 Continuous SLN
-This section introduces Continuous SLN, a generalization of SLN.
-Definition 4.1 (Continuous SLN).
-A Continuous Skip List Network (Continuous SLN) $\mathcal{N}=(V,h,E)$ is an SLN (Definition 2.1) except the height function $h$ is defined over the real numbers, i.e., $h:V\longrightarrow\mathbb{R}$.
-We prove that continuous skip graphs support a functioning P2P network by implementing the three primitives: route, join and leave.
-Internal Variables.
-A node $u$ holds the variables:
+  • The cost of all paths that include $x$, which by Lemma 1 is given by the Equation 4.
+Since we minimize this total over all pivot choices $x\in[l,r]$, the recurrence holds. ∎
+OSSLN with Bounded Height: While OSSLNu optimizes the total routing cost, it offers no guarantees regarding individual node degrees or the average network degree.
+Consequently, a single node could theoretically accumulate a near-linear number of edges, rendering a physical network implementation infeasible.
+Since the degree of a node is upper-bounded by twice its height, we can control the maximum degree by enforcing a height limit, $h_{\max}$.
+For the discrete SLN, capping the height at $h_{\max}$ strictly ensures that the maximum node degree does not exceed $2\times h_{\max}$.
+Definition 8 (OSSLN with Bounded Height $h_{\max}$).
+An OSSLN with bounded height $h_{\max}$ $\mathcal{N_{H}}=(V,h,E)$ is the discrete SLN (1) that minimizes the communication cost given by Equation 1 such that the height assignment for any of its nodes is at most $h_{\max}$.
+In the OSSLN with bounded height, the height assignment is of the form $h:V\longrightarrow[1,h_{\max}]$, where $h_{\max}$ is a parameter, typically bounded by $O(\log{n})$ or $O(1)$.
+Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$.
+Let $nl$, $1\leq nl\leq l$, denote the leftmost node id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let $nr$, $r\leq nr\leq n$, denote the rightmost node id, such that $\forall v\in[r,nr],\,h_{v}\leq h$.
+Finally, let $\textsc{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all nodes $v\in[l,r]$.
+Theorem 2.
+If $x\in[l,r]$ is a candidate pivot node with height $h\in[1,h_{\max}]$, then the following recurrence relation holds $\forall(nl,l,x,r,nr)$, $1\leq nl\leq l\leq x\leq r\leq nr\leq n$:
+$\text{OSSLN}(l,r,nl,nr,h,W)=\min\Bigg\{\text{OSSLN}(l,r,l,r,h-1,W),$
+$\min_{x\in[l,r]}\Bigg[\text{OSSLN}(l,x-1,nl,nr,h,W)+$
+$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
+$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\Bigg]\Bigg\}.$
+Proof.
+The proof is by strong induction on the size of the interval $l,r]\subseteq[1,n]$:
+Note that, by definition, $nl=L(x,h)$ and $nr=R(x,h)$ for any possible $h$.
+Furthermore, by construction, for an interval induced by $l$, $r$, nlnl, nrnr with bounded height $h$, assigning a pivot vertex with height equal to h1h-1, we have $nl^{\prime}=l$ and $nr^{\prime}=r$.
+Base case $(l=r)$: In this case, we have that $(x=l)$ and the set of paths $x$ belongs to is given by Lemma 1 By setting $nl=L(x,h)$ and $nr=R(x,h)$ in the proof, the vertex contribution to the total cost is given by:
+$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)$ (4)
+Induction step: For $(h=1)$, we can disregard the first half of the recursion, where we assign the pivot vertex a height lower than $h$. The total contribution of all vertices to the cost becomes:
+$\sum_{x\in[l,r]}\left(\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\right),$
+which, from the base case, is precisely what the second half of the recurrence equation will return for any choice of the pivot vertex.
+Let $1\leq l<r\leq n$ and $h\in[1,h_{\max}]$.
+Assume that the recurrence equation is true for all $h^{\prime},l^{\prime},r^{\prime}$ such that $(h^{\prime}<h)~\wedge~(r^{\prime}-l^{\prime})<(r-l)$.
+From previous step also assume that the equation is true for $(r^{\prime}-l^{\prime})=(r-l)$ with height $(h^{\prime}<h)$.
+Since, when assigning the pivot vertex with height lower than $h$, by construction, we have that $(nl=l)$ and $(nr=r$).
+The minimum cost for the solution for any $nl$ and $nr$ is the minimum between keeping the height of the pivot, $(h_{x}=h)$, or decreasing it, $h_{x}\leq(h-1)$, which is given by the minimum between:
+$\min_{x\in[l,r]}\Bigg[\textsc{OSSLN}(l,x-1,nl,nr,h,W)+$
+$\textsc{OSSLN}(x+1,r,nl,nr,h,W)+$
+$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\Bigg]$
+and OSSLN(l,r,l,r,h1,W)OSSLN(l,r,l,r,h-1,W), which is precisely what we want to prove. ∎
+The complexity of the above dynamic programming algorithm is $\mathcal{O}(n^{5}\cdot h_{\max})$, if we pre-compute an auxiliary matrix of aggregate weights $W^{\prime}(x,y)=\sum_{z\leq y}W(x,z),\forall(x,y),1\leq x\leq y\leq n$.
+4 Continuous Skip-List Networks
+This section introduces a generalization of SLN called C-SLN.
+Definition 9 (Continuous Skip List Network (C-SLN)).
+A C-SLN $\mathcal{N}=(V,h,E)$ is an SLN (1) except the height function $h$ is defined over the real numbers, i.e., $h:V\rightarrow\mathbb{R}$.
+In this section, we show that C-SLN offer a set of properties similar to classic SLN with arguably simpler proofs. Next, we detail key procedures of C-SLN, namely join, leave and route.
+4.1 Procedures
+We first present the internal variables used in our procedures. In particular, a node $u$ holds the variables:
   • $u$: the unique id of the node
@@
 Route.
-The route algorithm (Algorithm 1) takes the destination node id $v$ and hops from node to node until it finds the destination node, returning $(v,True)$ or reaches a dead end in some node $w\neq v$, return $(w,False)$.
-1 // code running at node $u$
-2 if $u$ = $v$ then
-3   return ($u$, True)
-4   else if $u<v$ then
-5     $S\leftarrow\{x\in N_{u}|u<x\leq v\}$
-6     if $S=\emptyset$ then
-7       return ($u$, False)
-8       else
-9         $w\leftarrow\max S$
-10         return $w.\text{route}(v)$
-11
-12         else
-13           $S\leftarrow\{x\in N_{u}|v\leq x<u\}$
-14           if $S=\emptyset$ then
-15             return ($u$, False)
-16             else
-17               $w\leftarrow\min S$
-18               return $w.\text{route}(v)$
-19
-Algorithm 1 route($v$)
+The route algorithm (Procedure 1) takes the destination node id $v$ and hops from node to node until it finds the destination node, returning $(v,True)$ or reaches a dead end in some node $w\neq v$, return $(w,False)$.
+// code running at node $u$
+if $u$ = $v$ then
+ return ($u$, True)
+end if
+else if $u<v$ then
+ $S\leftarrow\{x\in N_{u}|u<x\leq v\}$
+ if $S=\emptyset$ then
+    return ($u$, False)
+   end if
+ else
+    $w\leftarrow\max S$
+    return $w.\text{route}(v)$
+   end if
+end if
+else
+ $S\leftarrow\{x\in N_{u}|v\leq x<u\}$
+ if $S=\emptyset$ then
+    return ($u$, False)
+   end if
+ else
+    $w\leftarrow\min S$
+    return $w.\text{route}(v)$
+   end if
+end if
+Procedure 1 route($v$)
 Join.
-The join algorithm (Algorithm 2) inserts the node inside the Continuous SLN based on its anchor local variable.
-The potential neighbor changes of the other nodes propagate through the network with the find_neighbors algorithm (Algorithm 3).
-1
-2 // code running at node $u$
-3 if $\text{anchor}<u$ then
-4   $S\leftarrow\{x\in N_{\text{anchor}_{u}}|\text{anchor}<x\}$
-5
-6   else
-7     $S\leftarrow\{x\in N_{\text{anchor}_{u}}|x<\text{anchor}\}$
-8
-9     if $S=\emptyset$ then
-10       $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
-11       Update $D_{u}$
-12       else
-13         if $\text{anchor}<u$ then
-14           $\text{succ}\leftarrow\min S$
-15
-16           else
-17             $\text{succ}\leftarrow\max S$
-18
-19             $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
-20             $N_{u}\leftarrow N_{u}\cup\text{succ}.\text{find\_neighbors}(u,h_{v},-\infty)$
-21             Update $D_{u}$
-22
-Algorithm 2 join()
-1
-2 // code running at node $u$
-3 if current_height = -$\infty$ then
-4   $\text{anchor}_{u}\leftarrow v$
-5
-6  $N\leftarrow\emptyset$
-7   if current_height $h_{u}$ then
-8     $N\leftarrow\{u\}$
-9     $N_{u}\leftarrow N_{u}\cup\{v\}$
-10     Update $D_{u}$
-11     if $u<v$ then
-12       $N\leftarrow\{x\in N_{u}|x\leq v\text{ or }(h_{v}<D_{u}[x]\text{ and }h_{v}<h_{u})\}$
-13       for $x\in N_{u}$ do
-14         if $x<u\text{ and }h_{u}\leq D_{u}[x]\leq h_{v}$ then
-15           $N\leftarrow N\cup x.\text{find\_neighbors}(v,h_{v},h_{u})$
-16
-17
-18
-19          else
-20             $N\leftarrow\{x\in N_{u}|v\leq x\text{ or }(h_{v}<D_{u}[x]\text{ and }h_{v}<h_{u})\}$
-21             for $x\in N_{u}$ do
-22               if $u<x\text{ and }h_{u}\leq D_{u}[x]\leq h_{v}$ then
-23                 $N\leftarrow N\cup x.\text{find\_neighbors}(v,h_{v},h_{u})$
-24
-25
-26                 return NN
-Algorithm 3 find_neighbors($v$, $h_{v}$, current_height)
+The join algorithm (Procedure 2) inserts the node inside the Continuous SLN based on its anchor local variable.
+The potential neighbor changes of the other nodes propagate through the network with the find_neighbors algorithm (Procedure 4, in this paper’s appendix).
+// code running at node $u$
+if $\text{anchor}<u$ then
+ $S\leftarrow\{x\in N_{\text{anchor}_{u}}|\text{anchor}<x\}$
+end if
+else
+ $S\leftarrow\{x\in N_{\text{anchor}_{u}}|x<\text{anchor}\}$
+end if
+if $S=\emptyset$ then
+ $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
+   Update $D_{u}$
+end if
+else
+ if $\text{anchor}<u$ then
+    $\text{succ}\leftarrow\min S$
+   end if
+ else
+    $\text{succ}\leftarrow\max S$
+   end if
+ $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
+ $N_{u}\leftarrow N_{u}\cup\text{succ}.\text{find\_neighbors}(u,h_{v},-\infty)$
+   Update $D_{u}$
+end if
+Procedure 2 join()
 Leave.
-The leave algorithm (Algorithm 4, in this paper’s appendix) shows how an already-inserted node can leave the network.
-The potential neighbor changes propagate through the network with the delete_neighbors algorithm (Algorithm 5, in this paper’s appendix).
-5 The Uniform SLN Protocol
-After discussing the case of static SLN in the previous sections, we now model randomized P2P protocols.
-This section introduces a randomized, Continuous SLN called Uniform where each node chooses its height by picking a random real number in $(0,1)$ uniformly at random.
-Definition 5.1 (Randomized Continuous SLN).
-A randomized Continuous SLN is a probabilistic distribution over Continuous SLN.
-Protocol 1 (Uniform).
-Uniform is a randomized, continuous SLN that sets its heights to a real number in $(0,1)$ chosen uniformly at random.
-The rest of this section presents the properties of Uniform with respect to routing path lengths and maximum degrees.
-We show that a Continuous SLN constructed with Uniform has an expected constant degree and an expected routing path length of $O(\log n)$.
-Also, we can rule out rare bad outcomes in practice: with high probability, the routing path length is $O(\log n)$ and the maximum degree is $O(\log n)$.
-Those results directly imply that the three primitive described in the previous section have a logarithmic message complexity with high probability.
-Interestingly, those results are quite general and hold for a vast family of continuous probability laws, namely, as soon as the probability that two heights are equal is zero.
-Theorem 5.2 (Expected Path Length).
-The routing path between any two nodes has length smaller than $2\ln n$ on expectation.
-Proof 5.3.
+The leave algorithm (Procedure 3, in this paper’s appendix) shows how an already-inserted node can leave the network.
+The potential neighbor changes propagate through the network with the delete_neighbors algorithm (Procedure 5, in this paper’s appendix).
+// code running at node $u$
+if $\text{anchor}<u$ then
+ $S\leftarrow\{x\in N_{\text{anchor}_{u}}|\text{anchor}<x\}$
+end if
+else
+ $S\leftarrow\{x\in N_{\text{anchor}_{u}}|x<\text{anchor}\}$
+end if
+if $S=\emptyset$ then
+ $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{delete\_node}(u,h_{v},-\infty)$
+   Update $D_{u}$
+end if
+else
+ if $\text{anchor}<u$ then
+    $\text{succ}\leftarrow\min S$
+   end if
+ else
+    $\text{succ}\leftarrow\max S$
+   end if
+ $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{delete\_node}(u,h_{v},-\infty)$
+ $N_{u}\leftarrow N_{u}\cup\text{succ}.\text{delete\_node}(u,h_{v},-\infty)$
+   Update $D_{u}$
+end if
+Procedure 3 leave()
+4.2 Uniform C-SLN
+We now present a simple C-SLN called Uniform whose heights are chosen uniformly at random in $[0,1]$, and show that Uniform has the performances of state-of-the-art overlay network algorithms as both its routing search path and maximum degree are logarithmic with high probability.
+The properties of Uniform will be used to derive our main result on LASLiN in the next section.
+Algorithm 1 (Uniform).
+Uniform is a C-SLN whose initialization procedure sets its heights to a real number in $[0,1]$ chosen uniformly at random.
+Theorem 3 (Expected Routing Path Length).
+For Uniform, the routing path between any two nodes has length smaller than $2\ln n$ on expectation.
+Proof.
 Let $u$ and $v$ be two nodes.
@@
 The claimed upper bound on the expected path length follows as the same reasoning can be done for $X_{\text{down}}$.
-Theorem 5.4 (Routing Path Length with High Probability).
-The routing path between any two nodes has length smaller than $2e\cdot\ln n$ with probability greater than $1-\frac{2}{n}$.
-Proof 5.5.
+∎
+Theorem 4 (Routing Path Length with High Probability).
+For Uniform, the routing path between any two nodes has length smaller than $2e\cdot\ln n$ with probability greater than $1-\frac{2}{n}$.
+Proof.
 Let $u$ and $v$ be two nodes such that $u<v$, and let $X_{\text{up}}$ be the random variable equal to the increasing routing path length between $u$ and $v$ .
-Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to $1$ if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 5.2).
+Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to $1$ if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 3).
 We first show that all $1_{w}$ are independent random variables. Let $w$ and $w^{\prime}$ be two nodes such that $u<w<w^{\prime}$, it holds:
-$P[1_{w^{\prime}}=1\mid 1_{w}=1]$ $=P[h_{x}<h_{w^{\prime}}\forall x\in[w+1,w^{\prime}-1]\cap\max_{y\in[u,w]}h_{y}<h_{w}^{\prime}]$
-$=P[h_{y}<h_{w}^{\prime}\forall y\in[u,w^{\prime}-1]]$
-$=P[1_{w^{\prime}}=1]$
+$P[1_{w^{\prime}}=1\mid 1_{w}=1]$ $=$
+$P[h_{x}<h_{w^{\prime}}\forall x\in[w+1,w^{\prime}-1]\cap\max_{y\in[u,w]}h_{y}<h_{w}^{\prime}]$ $=$
+$P[h_{y}<h_{w}^{\prime}\forall y\in[u,w^{\prime}-1]]$ $=$
+$P[1_{w^{\prime}}=1]$
 Now that we showed independence, we use the following Chernoff’s bound: For $Y$ the sum of independent Bernoulli random variables (a Poisson trial) and $\mu=\mathbb{E}[Y]$, $\forall\delta>0$ it holds that:
@@
 $\leq P(X_{\text{up}}\geq e\cdot\ln n)+P(X_{\text{down}}\geq e\cdot\ln n)$
-$<\frac{2}{n}$
-Theorem 5.6 (Expected Degree).
-The expected degree of any node is lower than $4$.
-Proof 5.7.
+$<\frac{2}{n}\qed$
+Theorem 5 (Expected Degree).
+For Uniform, the expected degree of any node is lower than $4$.
+Proof.
 If $u$ and $w$ are two nodes, we define $N_{u}$ the random variable equal to the number of neighbors of $u$, and $1_{u,w}$ the indicator random variable equal to one if $u$ and $w$ are neighbors.
@@
 $\leq 4\cdot\sum_{i=1}^{n/2}\frac{1}{i\cdot(i+1)}=4\cdot\sum_{i=1}^{n/2}\frac{1}{i}-\frac{1}{i+1}$
-$=4\cdot(1-\frac{1}{n/2+1})\leq 4$
-Theorem 5.8 (Maximum Degree with High Probability).
-The maximum degree is lower than $4\ln n$ with probability greater than $1-\frac{1}{n^{3}}$.
-Proof 5.9.
+$=4\cdot(1-\frac{1}{n/2+1})\leq 4\qed$
+Theorem 6 (Maximum Degree with High Probability).
+For Uniform, the maximum degree is lower than $4\ln n$ with probability greater than $1-\frac{1}{n^{3}}$.
+Proof.
 Let $u$, $w$ and $w^{\prime}$ be three different nodes.
 We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to $1$ if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively.
-We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 5.4.
+We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 4.
 We first establish that each node taken individually has a degree lower than $\ln n$ with probability greater than $1-\frac{1}{n^{4}}$.
-Let $N_{u}$ be the random variable equal to the number of neighbors of $u$, we use the Chernov bound for Poisson trials used in the proof of Theorem 5.4:
+Let $N_{u}$ be the random variable equal to the number of neighbors of $u$, we use the Chernoff bound for Poisson trials used in the proof of Theorem 4:
 $P[N_{u}\geq 4\ln n]$ $<\left(\frac{e^{(\ln n)-1}}{(\ln n)^{\ln n}}\right)^{4}=\frac{(n/e)^{4}}{n^{4\ln\ln n}}<\frac{1}{n^{4}}$
@@
 We conclude with a union bound: the probability that at least one node has a degree greater than $4\ln n$ is upper bounded by $\sum_{v\in V}P[N_{v}\geq 4\ln n]\leq\frac{1}{n^{3}}$ and the claim holds.
-6 Learning-Augmented Continuous SLN
-In this section, we present LASLiN, a learning-augmented Continuous SLN protocol where each node can obtain a prediction of the future demand matrix at any time.
-We first show that, for general demands, any randomized, demand-oblivious P2P protocol can be off by a factor of $O(\log n/\log\log n)$ compared to an optimum static SLN.
+∎
+5 Learning-Augmented C-SLN
+In this section, we present LASLiN, a learning-augmented C-SLN algorithm, where each node can obtain a prediction of the future demand matrix at any time.
+We first show that, for general demands, any demand-oblivious overlay algorithm can be off by a factor of $O(\log n/\log\log n)$ compared to an optimum static SLN.
 This result provides a comparison point around which the consistency and robustness metrics of demand-aware algorithms can be evaluated.
-Theorem 6.1.
-We consider any randomized P2P protocol (Definition 5.1) that is ideal (Definition 2.4) and not learning-augmented (Definition 2.7).
-Then, there exists a demand matrix $W$ such that the considered protocol has a cost of $\Omega(\log n/\log\log n)\cdot|W|$ in expectation while an optimal static SLN protocol pays $|W|$, where $|W|$ is the total sum of the elements of $W$.
-Proof 6.2.
-As the considered protocol is ideal, there exists a constant $c$ such that the generated network has a maximum degree lower than $O(\ln^{c}n)$ with probability greater than 1O(1/n)1-O(1/n).
+Theorem 7.
+We consider any overlay algorithm that is not learning-augmented (6).
+Then, there exists a demand matrix $W$ such that the considered algorithm has a cost of $\Omega(\log n/\log\log n)\cdot|W|$ in expectation while an optimal static SLN algorithm pays $|W|$, where $|W|$ is the total sum of the elements of $W$.
+Proof.
+As the considered algorithms has a logarithmic maximum degree with high probability, there exists a constant $c$ such that the generated network has a maximum degree lower than $O(\ln^{c}n)$ with probability greater than 1O(1/n)1-O(1/n).
 Let $u$ be a node.
@@
 We consider the demand matrix that is zero everywhere except at the coefficient with row $u$ and column $v$.
-An optimal SLN protocol can clearly pay a cost of $1$, the considered randomized SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
+An optimal SLN algorithm can clearly pay a cost of $1$, the considered SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
 Finally, we cover the cases where the maximum degree is higher than $\ln^{c}n$: That case only contributes with an additive constant to the distance between $u$ and $v$ as a distance is lower than $n$.
 We considered all the cases and the claim follows.
-Protocol 2 (LASLiN).
-LASLiN is a randomized, continuous SLN that sets its heights by adding the input, integer prediction in $1,\lceil\ln n\rceil]$, and a real number in $(0,1)$ chosen uniformly at random.
-Theorem 6.3.
-LASLiN is a randomized, Continuous SLN that is $O(1)$-consistent and $O(\log^{2}n)$-robust.
-Proof 6.4.
-We start by proving that the maximum degree in LASLiN is lower than $4\ln^{2}n$ with high probability.
-Let $u$ be a node, let $p_{u}\in[1,\lceil\ln n\rceil]$ be the predicted height of $u$ and let $r_{u}\in(0,1)$ be the uniform random noise added on top of $p_{u}$.
-We partition the neighbors of $u$ into $p_{u}$ subsets called $S(1),S(2)\dots S(p_{u}+1)$: For all $i\in[1,p_{u}]$, $S(i)$ contains the neighbors with height between $i$ and $i+1$, the last set $S(p_{u}+1)$ contains the neighbors with height greater than $p_{u}+1$.
-According to Theorem 5.8, each set from $S(1)$ to $S(p_{u})$ contains fewer than $4\ln n$ nodes with probability greater than $1-\frac{1}{n^{4}}$ (this is the result before the last paragraph in the proof) — the set $S(p_{u}+1)$ contains at most two nodes.
-We conclude with a union bound that $u$ has fewer than $4\ln^{2}n$ with probability greater than $1-\frac{1}{n^{3}}$, and that the maximum degree is lower than $4\ln^{2}n$ with probability greater than $1-\frac{1}{n^{2}}$ also with a union bound.
-The $O(1)$-consistency is direct as the comparative optimal static SLN has its heights within a $O(\log n)$ factor without loss of generality as its maximum degree is bounded by $O(\log n)$ by assumption.
+∎
+Algorithm 2 (LASLiN).
+LASLiN is a learning-augmented C-SLN that takes an integer in $1,\lceil\ln n\rceil]$ as a prediction.
+Its initialization procedure set its height by adding the input prediction with a real number in ]0,1[]0,1[ chosen uniformly at random.
+Theorem 8.
+LASLiN is a C-SLN that is $O(1)$-consistent and $O(\log^{2}n)$-robust, while its maximum degree is $O(\log^{2}n)$ with high probability.
+Proof.
+We start by proving $O(1)$-consistency.
+Adding values in ]0,1[]0,1[ to the node’s heights does not remove the edges of the initial SLN.
+Let $u$ and $v$ be two nodes and consider the routing path $P_{u,v}=u,x_{1},x_{2}\dots v$ from $u$ to $v$ in the initial SLN.
+Now calling $P_{u,v}^{\prime}$ the path from $u$ to $v$ in the newly created C-SLN, we have that $P_{u,v}^{\prime}\subseteq P_{u,v}$; this can be show with a simple induction on the path length starting from $u$.
+All paths of the C-SLN have equal or shorted length than in the initial SLN with perfect predictions and $O(1)$-consistency follows.
 We now prove the robustness claim.
 A routing path between any two nodes $u$ and $v$ has two phases: first ascending (the heights always increase) and then descending (the heights always decrease).
-For both phase, the number of different encountered integer heights is upper-bounded by $O(\log n)$.
-For each encountered integer height, the expected routing path length is upper-bounded by $O(\log n)$ (Theorem 5.2).
-We finish the proof with an upper bound of each distance in the total weighted path lengths: $\sum_{u,v\in V}W(u,v)\cdot d(u,v)\leq O(\log^{2}n)\cdot|W|$ and the optimal static SLN pays at least $|W|$.
-We now define the notion of super-sparse matrix, which indicates a concentration of the weights of the matrix on a very small (logarithmic) number of coefficients.
-Definition 6.5 (super-sparse matrix).
-A matrix $W\in M_{n}(\mathbb{R})$ is super-sparse if there exists a constant $c\geq 1$ and a restricted subset of node pairs $S\subseteq V^{2}$ such that:
-$|S|$ $\leq O(\log^{c}n)\qquad\text{ and }$
-$\sum_{(u,v)\in S}W(u,v)$ $\geq\left(1-O(1/\log n)\right)\cdot|W|$
-Theorem 6.6.
-Assume that the demand matrix $W$ is super-sparse.
-We consider the protocol Binary LASLiN (B-LASLiN) where the input prediction $p$ can only have two values, $1$ or $2$.
-B-LASLiN is $O(\log\log n)$-consistent and $O(\log n)$-robust.
-In asymptotics, the performances are therefore always better than any demand-oblivious protocol, with an exponential improvement in case of perfect predictions.
-Proof 6.7.
-We start by proving that the maximum degree of each node is lower than $O(\log n)$ with high probability.
-As seen in the proof of Theorem 6.3, each node $u$ has fewer than $p_{u}\cdot\ln n$ neighbors with high probability, the claim follows since $p_{u}\leq 2$.
-We prove that the expected total weighted path length is lower than $O(\log n)\cdot|W|$ regardless of the input predictions.
-The proof is similar to the proof of Theorem 6.3 except, as the predictions are upper-bounded by a constant, we avoid the additional $\log n$ factor.
-Finally, we prove that the expected total weighted path length is lower than $O(\log\log n)\cdot|W|$.
-We consider the following set of predictions: let $S$ be the node pairs from Definition 6.5, we give prediction $2$ to the nodes that have at least one coefficient from $S$ in their row or column in $W$, otherwise $1$.
-We call $V_{2}$ ($V_{1}$) the set of nodes that have prediction 22 (11).
-As $W$ is super sparse, it holds that $|V_{2}|\leq O(\log^{c}n)$.
-$\sum_{u,v}W(u,v)\cdot d(u,v)$ $=\sum_{(u,v)\in S}W(u,v)\cdot d(u,v)+\sum_{(u,v)\in V^{2}\setminus S}W(u,v)\cdot d(u,v)$
-$\leq\sum_{u,v\in V_{2}}W(u,v)\cdot d(u,v)+\sum_{(u,v)\in V^{2}\setminus S}W(u,v)\cdot d(u,v)$
-$\leq O(\log\log n)\cdot|W|+O\left(\frac{1}{\log n}\right)\cdot|W|\cdot\ln n$
-$=O(\log\log n)\cdot|W|$
-To go from line 3 to 4 we used (1) Theorem 5.2 on the routing paths between the nodes of $V_{2}$ — there are $O(polylog(n))$ of those nodes, they have the largest heights of the SLN, so a path has expected length $\ln O(polylog(n))=O(\log\log n)$ — and on the routing paths from nodes in $V^{2}\setminus S$, and (2) we used Definition 6.5 to upper-bound the sum of coefficients in $W$ outside of $S$.
-Figure 6: Comparison of the costs of multiple P2P networks against different demand matrix distributions.
-The demand matrices are generated based on the Zipfian distribution with varying skewness parameter $\zeta\in\{1.1,1.2\dots 2.0\}$, each coefficient is drawn independently and identically distributed.
-For each demand matrix distribution, we compare the cost of three P2P networks: the optimal static SLN, LASLiN, and the demand-oblivious Uniform SLN.
+For both phase in the initial SLN, the number of different encountered integer heights is upper-bounded by $O(\log n)$.
+For each encountered integer height, the expected routing path length in the C-SLN is upper-bounded by $O(\log n)$ (Theorem 3) with high probability, we deduce that the routing path is lower than $O(\log^{2}n)$ with high probability using a union bound on the integer heights of the initial SLN.
+We now use this upper bound of each distance in the total weighted path lengths: $\sum_{u,v\in V}W(u,v)\cdot d(u,v)\leq O(\log^{2}n)\cdot|W|$ and the optimal static SLN pays at least $|W|$.
+Finally, we show the claim on the maximum degree.
+Let $u$ be a node.
+Fixing an integer $i\in[0,p_{u}]$, in the newly created C-SLN $u$ has fewer than $\ln n$ neighbors with initial integer heights $i$ with high probability (Theorem 6).
+We complete the proof with a union bound on all possible $O(\log n)$ heights of the initial SLN.
+∎
+Figure 3: Consistency (Pareto $\alpha\in[0.2,2]$, $n=64$, $\eta=0$).
+(a) $\alpha=0.2$
+(b) $\alpha=0.4$
+(c) $\alpha=0.6$
+(d) $\alpha=0.8$
+(e) $\alpha=1.0$
+(f) $\alpha=1.2$
+(g) $\alpha=1.4$
+(h) $\alpha=1.6$
+(i) $\alpha=1.8$
+(j) $\alpha=2.0$
+Figure 4: Smoothness against Stale Noise (Pareto demand, $n=64$): average communication cost relative to uOPT.
+(a) $\alpha=0.2$
+(b) $\alpha=0.4$
+(c) $\alpha=0.6$
+(d) $\alpha=0.8$
+(e) $\alpha=1.0$
+(f) $\alpha=1.2$
+(g) $\alpha=1.4$
+(h) $\alpha=1.6$
+(i) $\alpha=1.8$
+(j) $\alpha=2.0$
+Figure 5: Smoothness against Adversarial Noise (Pareto demand, $n=64$): average communication cost, relative to uOPT.
+(a) pFabric traces ($\eta=0$)
+(b) pFabric trace_0_1\_0\_1
+(c) pFabric trace_0_5\_0\_5
+(d) pFabric trace_0_8\_0\_8
+Figure 6: pFabric demand under Adversarial Noise ($n=144$): consistency (a) and smoothness (b),(c),(d).
+(a) Pareto $\alpha=0.2,n=64$
+(b) pFabric trace_0_5,n=144\_0\_5,n=144
+Figure 7: Node degree CDF ($\eta=0$): the vertical lines indicate the average node degree.
 6 Experimental Evaluation
-Demand matrices
-In our experiments, we generate communication demand based on the Zipfian distribution [1], which receives a parameter $\zeta>0$, that indicates the distribution’s skewness.
-Roughly speaking, lower values of $\zeta$ result in a sparser demand matrix.
-The number of nodes in all experiments was set to $n=32$, and each experiment was repeated $30$ times.
-Noise generation
-As common in the literature [20], we study the robustness of our algorithms by introducing noise in the predicted data.
-The most used type of noise is additive, often gaussian [26] or lognormal [20].
-This type of noise however almost did not impact the performance of our algorithms.
-Instead, we test two other models: the demand error recomputes a portion (x-axis) of the coefficients with the same Zipfian distribution, and the unit heights enforce that each node can receive a predicted height of $1$ with a certain probability (x-axis).
-Results
-In Figure 6 we evaluate the consistency of LASLiN in scenarios where the demand follows the Zipfian distribution, with varying parameter $\zeta\in[1.1,2]$, and where the prediction advice is perfect (i.e., there is zero noise in the demand matrix).
-We compare the total communication cost (as defined in (2.5)) of LASLiN to that of the static optimum SLN and the demand-oblivious SLN (Uniform) baselines.
-We can see that, in contrast to the demand-oblivious SLN, LASLiN obtains close-to-optimum total communication cost in all workloads, regardless of their spatial locality, i.e., LASLiN is consistent.
-In Figure 7 we evaluate the robustness of LASLiN in scenarios with increasing demand noise (x-axis).
-We compare the total cost of LASLiN to that of the static optimum SLN, the demand-oblivious (Uniform) SLN, and the SLN baselines.
-We can observe that the SLN performs increasingly worse (has higher cost) the greater the amount of change (noise) in the input demand matrix.
-LASLiN, on the other hand, achieves consistently lower communication cost (than the former), regardless of the type of noise and regardless of the prediction fault probability, i.e., LASLiN is robust relative to both type and magnitude of prediction errors.
-Figure 7: Cost of different P2P networks under varying noise amplitudes on the demand matrix predictions. Simulation results: robustness ($\zeta=1.2$)
+In this section, we present our experimental setup and results.
+Algorithms and baselines.
+To evaluate the performance of LASLiN, we implemented the following algorithms:
+  • Sln: Demand-Oblivious Discrete SLN (1, with the height of each node having been assigned the number of consecutive independent trials in which a random variable $x\in[0,1]$ satisfied $x<0.5$);
+  • uOPT: Unbounded Optimum Static SLN (4) on predicted demand;
+  • bOPT: Bounded Optimum Static SLN (8 on predicted demand, with $h_{\max}=6$);
+  • bLasin: LASLiN (6, with bOPT as input prediction);
+Demand matrices.
+In our experiments, we used both synthetic and real-world workloads.
+The synthetic workloads were based on a Pareto distribution, with the shape parameter $\alpha\in[0.2,2.0]$ (the higher the value of $\alpha$, the closer the demand to a uniform distribution).
+For each value of $\alpha$, we generated $30$ distinct demand matrices on a network with $64$ nodes.
+To validate our algorithms on real-world data, we utilized three pFabric traces Alizadeh et al.
+(2013), generated by executing simulation scripts in NS2 on a network with $144$ nodes.
+Noise generation.
+In line with some previous related work  Lykouris and Vassilvitskii (2018), we evaluate the smoothness of our algorithms by introducing noise, or prediction error, either to the demand weight values (between pairs of nodes) or to the heights assigned to nodes by an SLN algorithm.
+In particular, we implemented two noise models:
+  • Stale Noise: This model aims to simulate the effect of an outdated input prediction through a probabilistic swapping mechanism.
+For every node pair in the demand matrix, the demand weight value is replaced by a different value, taken from a historical or reference workload, with probability $\eta$ (the noise intensity parameter).
+  • Adversarial Noise: This model simulates the effect of the presence of malicious nodes in the network.
+For every network node, the height value assigned to it by the Skip List algorithm is reset to $1$, with probability $\eta$.
+The Stale Noise model represents scenarios with delayed updates of input prediction values, caused by, e.g., network reconfigurations or synchronization errors.
+We restricted the evaluation of smoothness against stale noise to demand matrices based on Pareto distributions.
+[^1] We excluded pFabric traces from this part of the evaluation, as the dataset lacked sufficient independent traces to generate reliable reference models.
+Given a workload $W_{i},1\leq i<30$, we designated the subsequent workload, $W_{i+1}$, as the reference source for the current workload, $W_{i}$.
+Specifically, for each communication pair $(x,y)$, the actual demand weight $W_{i}(x,y)$ was replaced by the reference weight $W_{i+1}(x,y)$, with probability $\eta$.
+In the Adversarial Noise model, adversarial agents reset their own node heights to a baseline value of $1$, with probability $\eta$, eliminating structural (SLN) shortcuts and increasing the total routing work (communication cost).
+The objective of this experiment is to demonstrate the efficacy of a simple countermeasure: by increasing each node’s height value with a small randomized value, we can effectively safeguard the network’s performance against such adversarial intent.
+Metrics.
+We used the following performance metrics:
+  • Communication cost (3);
+  • Consistency (7) and smoothness;
+  • Node degree cumulative distribution function (CDF): percentage of nodes with degree $\leq x,x\in[1,n)$, and the average node degree.
+Results.
+In Figures 3 and 6(a), we evaluate the consistency of LASLiN under Pareto-distributed demand with variable shape parameter ($\alpha\in[0.2,2.0]$), assuming perfect prediction advice ($\eta=0$).
+We compare the (average) total communication cost of LASLiN (bLaslin) against three baselines: the bounded static optimum (bOPT), the unbounded static optimum (uOPT), and the demand-oblivious discrete SLN (Sln).
+The results demonstrate that, unlike the demand-oblivious baseline, LASLiN achieves a total communication cost close to the optimum across all simulated workloads, regardless of spatial locality ($\alpha$).
+This empirically confirms that LASLiN is consistent.
+Notably, LASLiN, despite being projected onto the bOPT structure (bLaslin), i.e., using the solution of bOPT as its input prediction, actually achieves better communication cost than bOPT, even in noiseless scenarios (Fig.
+3).
+This performance gain is enabled by the slight increase in the network’s average and maximum degrees, as shown in Figures 7(a) and 7(b).
+This behavior has been expected, since bLaslin effectively breaks height ties, inherent in the bOPT solution.
+By adding a small random remainder to the predicted heights, nodes that have previously shared the same rank (height) become differentiated, generating additional (SLN) shortcut links that further reduce routing cost.
+In Figure 4, we evaluate the smoothness of LASLiN against increasing levels of stale noise.
+Both uOPT and bOPT are significantly impacted by the intensity of the stale noise (increasing $\eta$ on the x-axis), especially for lower $\alpha$ values, where the workload is more skewed.
+In contrast, bLaslin consistently achieves lower communication cost than its input prediction (bOPT).
+This demonstrates that LASLiN remains robust relative to both the type and magnitude of input prediction errors.
+Additionally, we observe that, as $\alpha$ increases, approaching a uniform demand distribution, the performance of the optimum SLN converges to that of the discrete demand-oblivious SLN.
+The strongest advantage of LASLiN is observed in the presence of adversarial noise (Figures 5 and 6(b,c,d), note that the y-axis of these plots is shown in the log scale).
+By simply projecting the LASLiN mechanism onto the predicted node heights, we successfully mitigate the disruption that could be caused by adversarial agents.
+We observe that while both uOPT and bOPT degrade progressively as noise intensity increases, eventually underperforming even the demand-oblivious discrete SLN baseline, bLaslin consistently achieves lower communication costs than both its reference prediction (bOPT) and the lower bound baseline (uOPT).
+Finally, it is important to note that the average and the maximum node degree values obtained in the simulations (Fig.
+7) are in alignment with the theoretical expectation (Theorem 5, Theorem 6 ).
+In particular, for the discrete SLN, the maximum degree is approximately $2\cdot\log n$ (around $12$ for $n=64$), which, in the bounded case, was exactly equal to the limit of $2\cdot h_{\max}=12$.
+Notably, the average degree remained tightly consistent across all algorithms.
+This confirms that the degree distribution of an SLN is heavily skewed, with the vast majority of nodes maintaining a low degree.
 7 Related Work
-Improved designs of P2P networks.
-P2P networks have been the focus of computer scientists for a long while (see [19] for a detailed survey).
-In terms of P2P networks with constant degree, the notable contribution of [23] that extends the work of [13], giving a general P2P construction such that for a network with degree $d$ and $n$ nodes, you can have guaranteed path length of $\Theta(\log_{d}n)$.
-Lately, improving the design of networks given prior information about communication demand has grabbed the attention of researchers.
-In particular, the work of [7] set the scene for demand-aware network design, which was then explored more with particular focus on datacenter networking [11, 24, 12].
-Works closest to ours are by [6, 10] in which the authors discuss a self-adjusting variant of skip list network: one that adjusts over time based on the incoming demand, which is in contrast with our work that focuses on an optimal skip list network with static height given the communication demand as the input.
-This is in particular important, as self-adjustments can introduce additional cost (e.g., in terms of delay) to the system.
-Also, the work of [21] discussed how a skip-list construction can be used for peer-to-peer communications, however they do not provide any theoretical upper bound for the communication time.
-Lastly, we point out to other extensions of skip list, namely Skip Graph [3] and SkipNet [16] that, to facilitate P2P communications, consider expected degree to be $O(\log n)$, in contrast to our design, which has a constant degree in expectation.
+Improved designs of overlay networks.
+Overlay networks have been the focus of computer scientists for a long while (see Lua et al.
+(2005) for a detailed survey).
+Many overlay networks of constant epected degree and logarithmic paths lengths are known, for example Naor and Wieder (2007) that extends the work of Fraigniaud and Gauron (2006), giving a general overlay construction such that for a network with degree $d$ and $n$ nodes, it is possible to guarantee path length of $\Theta(\log_{d}n)$.
+Lately, improving the design of networks given prior information about communication demand has attracted the attention of researchers.
+In particular, the work of Avin and Schmid (2018) set the scene for demand-aware network design, which was then explored more with particular focus on datacenter networking de Oliviera Souza et al.
+(2022); Pourdamghani et al.
+(2023); Figiel et al.
+(2025).
+Works closest to ours are by Avin et al.
+(2020c); Ciriani et al.
+(2007) in which the authors discuss a self-adjusting variant of skip list network: one that adjusts over time based on the incoming demand, which is in contrast with our work that focuses on an optimal skip list network with static height given the communication demand as the input.
+This is particularly important, as self-adjustments can introduce additional cost (e.g., in terms of delay) to the system.
+Also, the work of Mandal et al.
+(2015) discussed how a skip-list construction can be used for peer-to-peer communications, however they do not provide theoretical upper bounds for the cost of their overlay network.
+Lastly, we point out to other extensions of skip list, namely Skip Graph Aspnes and Shah (2003) and SkipNet Harvey et al.
+(2003) that, to facilitate overlay communications, consider expected degree to be $O(\log n)$, in contrast to our design, which has a constant degree in expectation.
 Learning-augmented algorithms.
-Given the rise of machine learning algorithms (that are prone to inherent unbounded failures), and after the pioneering work of [26] on ski rental & job scheduling, the focus has been shifted on studying consistency and robustness of algorithms augmented with the learned advice.
-Such an approach has been used in various applications so far222see a full list of recent works in the following website: algorithms-with-predictions.github.io, for example caching [20], online TSP [15], list update [8].
+Given the rise of machine learning algorithms (which inherently comes with some errors), and after the pioneering work of Purohit et al.
+(2018) on ski rental & job scheduling, the focus has been shifted on studying consistency and robustness of algorithms augmented with the learned advice.
+Such an approach has been used in various applications so far [^2] See a full list of recent works in the following website: algorithms-with-predictions.github.io, for example caching Lykouris and Vassilvitskii (2018), online TSP Gouleakis et al.
+(2023), list update Azar et al.
+(2025).
 Some of the more recent works focus on improving single-source search data structures.
-Initially, the works of [18, 9] provided learning-augmented search trees.
-The works that are closest to our work are by [14, 30], which studied learning-augmented single-source skip lists.
+Initially, the works of Lin et al.
+(2022); Chen et al.
+(2025) provided learning-augmented search trees.
+The works that are closest to our work are by Fu et al.
+(2025); Zeynali et al.
+(2024), which studied learning-augmented single-source skip lists.
 In contrast, we consider a more challenging model, where the focus is on pairwise communications rather than communication from a single source.
-Recently, we became aware of a not peer-reviewed manuscript [2] that aims to augment P2P networks with learned advice; however, this paper looks into improving self-stabilization and recovery of P2P networks, not improving communication delay.
-9 Conclusion
-In this work, introduced a continuous variant of Skip List Network (SLN), Continuous SLN.
-This variant is arguably easier to manipulate and analyze.
-In addition we show that we can use it to reach strong robustness properties even in an adversarial, fully distributed setting.
-Lastly, through experimental evaluation, we observed that our suggested algorithms outperform the theoretical results.
-As a future work, one can think of an alternative algorithm for OSSLN with a lower computational complexity.
+Recently, we became aware of a not peer-reviewed manuscript Aradhya and Scheideler (2025) that aims to augment overlay networks with learned advice; however, this paper looks into improving self-stabilization and recovery of overlay networks, not improving communication delay.

### v3

@@
 As our third contribution, we present a learning-augmented overlay algorithm, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform algorithm with a predicted SLN to strike consistency and robustness guarantees (7).
-For each node, LASLiN takes as prediction an integer between $1$ and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $[0,1]$ uniformly at random and adds it to the integer heights to obtain its final heights as a C-SLN; it may then join the network using the join() procedure.
+For each node, LASLiN takes as prediction an integer between 11 and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $[0,1]$ uniformly at random and adds it to the integer heights to obtain its final heights as a C-SLN; it may then join the network using the join() procedure.
 We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $O(\log^{2}n)$ factor away from the optimum static SLN in case of adversarial predictions ($O(\log^{2}n)$-robust).
 0123412345678
-Figure 1: A SLN with eight nodes. The node ids range from $1$ to $8$. The double-arrowed, horizontal lines are the edges of the SLN.
+Figure 1: A SLN with eight nodes. The node ids range from 11 to 88. The double-arrowed, horizontal lines are the edges of the SLN.
 0123412345678
@@
 Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$.
-Let $nl$, $1\leq nl\leq l$, denote the leftmost node id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let $nr$, $r\leq nr\leq n$, denote the rightmost node id, such that $\forall v\in[r,nr],\,h_{v}\leq h$.
+Let nlnl, $1\leq nl\leq l$, denote the leftmost node id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let nrnr, $r\leq nr\leq n$, denote the rightmost node id, such that $\forall v\in[r,nr],\,h_{v}\leq h$.
 Finally, let $\textsc{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all nodes $v\in[l,r]$.
@@
 Since, when assigning the pivot vertex with height lower than $h$, by construction, we have that $(nl=l)$ and $(nr=r$).
-The minimum cost for the solution for any $nl$ and $nr$ is the minimum between keeping the height of the pivot, $(h_{x}=h)$, or decreasing it, $h_{x}\leq(h-1)$, which is given by the minimum between:
+The minimum cost for the solution for any nlnl and nrnr is the minimum between keeping the height of the pivot, $(h_{x}=h)$, or decreasing it, $h_{x}\leq(h-1)$, which is given by the minimum between:
 $\min_{x\in[l,r]}\Bigg[\textsc{OSSLN}(l,x-1,nl,nr,h,W)+$
@@
 We will give an upper bound of $X_{\text{up}}$, the same upper bound will hold for $X_{\text{down}}$ by symmetry.
-Let $w$ be a node between $u$ and $v$ in the SLN, we note $1_{w}$ the indicator random variable equal to $1$ if (1) $w$ is part of the path between $u$ and $v$, and (2) if the height increases when hoping to $w$.
+Let $w$ be a node between $u$ and $v$ in the SLN, we note $1_{w}$ the indicator random variable equal to 11 if (1) $w$ is part of the path between $u$ and $v$, and (2) if the height increases when hoping to $w$.
 It holds $X_{\text{up}}=\sum_{w\in[u+1,v]}1_{w}$.
@@
 Let $u$ and $v$ be two nodes such that $u<v$, and let $X_{\text{up}}$ be the random variable equal to the increasing routing path length between $u$ and $v$ .
-Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to $1$ if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 3).
+Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to 11 if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 3).
 We first show that all $1_{w}$ are independent random variables. Let $w$ and $w^{\prime}$ be two nodes such that $u<w<w^{\prime}$, it holds:
@@
 Theorem 5 (Expected Degree).
-For Uniform, the expected degree of any node is lower than $4$.
+For Uniform, the expected degree of any node is lower than 44.
 Proof.
@@
 Let $u$, $w$ and $w^{\prime}$ be three different nodes.
-We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to $1$ if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively.
+We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to 11 if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively.
 We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 4.
@@
 We consider the demand matrix that is zero everywhere except at the coefficient with row $u$ and column $v$.
-An optimal SLN algorithm can clearly pay a cost of $1$, the considered SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
+An optimal SLN algorithm can clearly pay a cost of 11, the considered SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
 Finally, we cover the cases where the maximum degree is higher than $\ln^{c}n$: That case only contributes with an additive constant to the distance between $u$ and $v$ as a distance is lower than $n$.
@@
   • Adversarial Noise: This model simulates the effect of the presence of malicious nodes in the network.
-For every network node, the height value assigned to it by the Skip List algorithm is reset to $1$, with probability $\eta$.
+For every network node, the height value assigned to it by the Skip List algorithm is reset to 11, with probability $\eta$.
 The Stale Noise model represents scenarios with delayed updates of input prediction values, caused by, e.g., network reconfigurations or synchronization errors.
@@
 Specifically, for each communication pair $(x,y)$, the actual demand weight $W_{i}(x,y)$ was replaced by the reference weight $W_{i+1}(x,y)$, with probability $\eta$.
-In the Adversarial Noise model, adversarial agents reset their own node heights to a baseline value of $1$, with probability $\eta$, eliminating structural (SLN) shortcuts and increasing the total routing work (communication cost).
+In the Adversarial Noise model, adversarial agents reset their own node heights to a baseline value of 11, with probability $\eta$, eliminating structural (SLN) shortcuts and increasing the total routing work (communication cost).
 The objective of this experiment is to demonstrate the efficacy of a simple countermeasure: by increasing each node’s height value with a small randomized value, we can effectively safeguard the network’s performance against such adversarial intent.